### v3

# Revisiting Randomization in Greedy Model Search

Abstract
Feature subsampling is a core component of random forests and other ensemble methods. While recent theory suggests that this randomization acts solely as a variance reduction mechanism analogous to ridge regularization, these results largely rely on base learners optimized via ordinary least squares. We investigate the effects of feature subsampling on greedy forward selection, a model that better captures the adaptive nature of decision trees. Assuming an orthogonal design, we prove that ensembling with feature subsampling can reduce both bias and variance, contrasting with the pure variance reduction of convex base learners. More precisely, we show that both the training error and degrees of freedom can be non-monotonic in the subsampling rate, breaking the analogy with standard shrinkage methods like the lasso or ridge regression. Furthermore, we characterize the exact asymptotic behavior of the estimator, showing that it adaptively reweights OLS coefficients based on their rank, with weights that are well-approximated by a logistic function. These results elucidate the distinct role of algorithmic randomization when interleaved with greedy optimization.
Keywords: Ensemble Learning; Forward Selection; Greedy Model Search; Random Forest; Regression
1 Introduction
Introduced by breiman2001random, random forests (RFs) are ensembles of decision trees, each grown via greedy, recursive partitioning of the input space (breiman1984classification). Each tree is independently randomized in two ways. First, it is trained on a bootstrap sample of the data. Second, at every internal node it considers only a randomly drawn subset of $m$ (out of $p$ total) features when searching for the next split.
This combination of aggregation and randomized tree construction has proven remarkably effective, making RFs among the most widely used machine learning methods today. Their performance is often close to state-of-the-art on tabular data (caruana2006empirical; fernandez2014we; olson2018data; grinsztajn2022tree), which is prevalent across healthcare, finance, and scientific domains.
There has been significant theoretical work on the consistency of RFs, especially in high dimensions (scornet2015consistency; klusowski2020sparse; syrgkanis2020estimation; chi2022asymptotic; klusowski2024large; mazumder2024convergence), but there is not yet a complete consensus on how randomization improves prediction performance. breiman2001random originally explained the benefits of randomization as increasing diversity among the trees, thereby reducing the variance in their predictions. Interestingly, in his concluding remarks, he speculated that randomization may “act to reduce bias” but that “the mechanism for this is not obvious”. These two explanations have had divergent fates, with hastie2009elements arguing in their influential textbook that “the improvements in prediction obtained by bagging or random forests are solely a result of variance reduction”.
Seminal work by mentch2020randomization elaborated on this “pure variance reduction” hypothesis by offering a quantitative analysis of the role played by the feature subsampling ratio $\gamma=m/p$. In their numerical experiments, they found that decreasing $\gamma$ decreases the degrees of freedom in RFs. To further explore the generality of this phenomenon, they applied ensembling with feature subsampling to the greedy forward selection (FS) algorithm for sparse linear regression (draper1966applied), finding that decreasing $\gamma$ also decreases the degrees of freedom of the resulting family of estimators. Finally, in concurrent work with lejeune2020implicit, they analyzed ensembling with feature subsampling on ordinary least squares (OLS) estimators. Such an ensemble remains a linear estimator, unlike ensembles of FS. Assuming orthonormal features, they found that the resulting regression coefficients were shrunk from the OLS values by a factor of $\gamma$, making the algorithm equivalent to ridge regression. They hence concluded that feature subsampling “serves as a form of implicit regularization” and that $\gamma$ “serves much the same purpose as the shrinkage penalty in explicitly regularized regression procedures like lasso and ridge regression”.
Recent work building on mentch2020randomization has proceeded in two directions. In one direction, patil2024asymptotically relaxed the orthonormal feature assumption to one they called “asymptotic freeness” (relative to the feature subsampling operator). Under this assumption, they showed that ensembling OLS estimators with column-wise sampling (“feature subsampling”) is asymptotically equivalent to ridge regression, albeit with inflated variance. Meanwhile, other works studying row-wise sampling (“data subsampling”) established a broad theoretical equivalence between this form of randomized ensembling and ridge penalization (du2023subsample; patil2023generalized). Together, these works seem to strengthen the analogy between various forms of subsampling randomization and ridge-like shrinkage.
In the other direction, studies by curth2024random, liu2025randomization, mei2024exogenous, and revelas2025random suggest that feature subsampling can also reduce bias in RFs via two mechanisms. The first is enlarging the effective model class (curth2024random). The second is overcoming failures of greedy optimization (liu2025randomization). These results challenge the “pure variance reduction” hypothesis of hastie2009elements and suggest that the shrinkage analogy in mentch2020randomization seems appropriate when describing ensembles of base learners fit with convex optimization, but this analogy becomes strained when subsampling is interleaved with greedy optimization, as is the case for RFs. Existing theory on subsampling-based ensembling primarily concerns non-adaptive or analytically tractable base learners arising from convex optimization, or oracle versions of RFs (mei2024exogenous). For settings in which feature subsampling is interleaved with greedy, adaptive fitting, evidence for bias-reduction mechanisms remains largely empirical.
The goal of our paper is to theoretically investigate the nuances of ensembling with feature subsampling that arise in this more complex scenario. Inspired by mentch2020randomization, we study their ensembling strategy as applied to FS rather than RFs. This model captures the key adaptive, greedy behavior of RFs and allows for rigorous mathematical analysis that would otherwise be intractable. We assume a fixed design setting, under which the expected training error captures the contribution of bias and the degrees of freedom quantify the estimator’s sensitivity to noise and thus its variance component. We further assume orthogonal features, in which case FS is equivalent to best subset selection (BSS, hastie2020extended). Letting FS($k$) and EFS($k,m$) refer to $k$-step forward selection and its ensemble, respectively, we obtain the following theoretical results.
  • For $p$ large enough, the training error of EFS($k,m$) is non-monotonic in $m$, with a minimum achieved at an intermediate value.
  • If the degrees of freedom of FS($k$) forms a concave sequence in $k$, then the degrees of freedom of EFS($k,m$) is increasing in $m$.
  • If the concavity condition does not hold, then the degrees of freedom of EFS($k,m$) may be non-monotonic in $m$, with a minimum achieved at an intermediate value.
These results provide rigorous proof that feature subsampling randomization can simultaneously reduce both bias and variance. On the other hand, variance reduction is not guaranteed but depends on certain structural assumptions. Furthermore, in both ridge and lasso regression, the expected training error (resp. degrees of freedom) is monotonically increasing (resp. monotonically decreasing) in the regularization parameter $\lambda$ (zou2007degrees; hastie2009elements), yielding an exact trade-off between the two quantities. Since this monotonic relationship does not hold for $m$ in EFS($k,m$), we see that shrinkage is not quite an appropriate analogy for the role of randomization in this setting.
In addition, focusing on the asymptotic regime $m,p\to\infty$ with $m/p\to\gamma$, we characterize the exact effect that $\gamma$ has on the regression coefficients of EFS($k,m$). Note that it is standard practice to take $m=\lfloor p/3\rfloor$ in RFs, corresponding to $\gamma=1/3$. Specifically, we show that EFS($k,m$) adaptively reweights OLS coefficients, creating a smooth approximation to the hard thresholding utilized by FS($k$). These weights are well-approximated by a logistic function in the OLS coefficient rankings, with inflection point roughly equal to $k$ and growth rate depending on $\gamma$.
Varying $\gamma$ (equivalently, $m$) for a fixed $k$ can be thought of as reallocating a fixed weight budget among the different features, thereby searching over a larger family of models compared to FS($k$). Since increasing $\gamma$ simultaneously shrinks some coefficients and inflates others, this reveals another difference from ridge and lasso regression, which shrink all coefficients towards zero. Indeed, $\gamma$ has more affinity with the $\alpha$ parameter in elastic net (zou2005regularization), which interpolates between the $\ell_{1}$ and $\ell_{2}$ penalties. However, the family of solutions obtained by EFS($k,m$) is not equivalent to that from elastic net, revealing yet more subtleties in the relationship between explicit regularization and “algorithmic regularization”.
We provide proof sketches for some of our main results. Full details appear in the Supplementary Material (chen2025revisiting-supp).
2 Preliminaries
2.1 Set-up and notation
Assume a generative model $y_{i}=f(\mathbf{x}_{i})+\varepsilon_{i}$, where $f(\mathbf{x})=\boldsymbol{\beta}^{\top}\mathbf{x}$ is a linear function and $\varepsilon_{1},\varepsilon_{2},\ldots,\varepsilon_{n}$ are IID noise variables with zero mean and variance $\sigma^{2}$. Strictly speaking, we do not need to assume a well-specified linear model for our main results to hold, but we do so here for conceptual simplicity. We assume $\mathbf{X}=[\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{n}]^{\top}\in\mathbb{R}^{n\times p}$ is a fixed design matrix. For vectors $\mathbf{a},\mathbf{b}\in\mathbb{R}^{n}$, we define the inner product as $\langle\mathbf{a},\mathbf{b}\rangle=\mathbf{a}^{\top}\mathbf{b}/n$, with associated norm $\|\mathbf{a}\|^{2}=\langle\mathbf{a},\mathbf{a}\rangle$. For each feature index $j\in[p]$, let $\mathbf{x}_{\cdot j,p}=(x_{1j},x_{2j},\dots,x_{nj})^{\top}$ denote the $j$-th feature vector, which we assume to be normalized so that $\|\mathbf{x}_{\cdot j,p}\|=1$. We use regular font to denote functions such as the coordinate functions $x_{j,p}$.
2.2 Algorithms
Forward selection.
Starting with an empty active set $\mathcal{M}_{0}=\emptyset$, forward selection (FS) proceeds for $l=1,2,\dots,k$ by updating $\mathcal{M}_{l-1}$ with the index
$j_{l}=\operatorname*{arg\,min}_{j\in\mathcal{V}_{l}}\|\mathbf{y}-\mathbf{P}_{\mathcal{M}_{l-1}\cup\{j\}}\mathbf{y}\|=\operatorname*{arg\,max}_{j\in\mathcal{V}_{l}}\frac{|\langle\mathbf{r}_{l-1},\mathbf{x}_{\cdot j,p}\rangle|}{\|\mathbf{P}_{\mathcal{M}_{l-1}}^{\perp}\mathbf{x}_{\cdot j,p}\|},$ (1)
where $\mathcal{V}_{l}=[p]\setminus\mathcal{M}_{l-1}$ is the set of eligible indices. Here, $\mathbf{P}_{\mathcal{M}_{l}}$ is the projection matrix onto the column span of the selected features, $\mathbf{P}_{\mathcal{M}_{l}}^{\perp}$ projects onto its orthogonal complement, and $\mathbf{r}_{l-1}=\mathbf{y}-\hat{\mathbf{f}}_{l-1,p}$ denotes the residual vector at step l1l-1. We set
$\hat{\mathbf{f}}_{l,p}=\mathbf{P}_{\mathcal{M}_{l}}\mathbf{y}.$
After $k$ steps, the resulting predictive function $\hat{f}_{k,p}$ is referred to as the FS($k$) estimator.
Ensemble forward selection.
Following mentch2020randomization, we introduce randomization into each FS($k$) step by modifying (1) so that the maximization is taken over a randomly drawn subset. Specifically, at step $l$, we sample a candidate set $\mathcal{V}_{l}\subseteq[p]\setminus\mathcal{M}_{l-1}$ of size $m$ uniformly without replacement. If $|[p]\setminus\mathcal{M}_{l-1}|<m$, we use all remaining features without randomization. The update index $j_{l}$ is then chosen as in (1), but restricted to $j\in\mathcal{V}_{l}$. Stopping after $k$ steps yields a randomized FS($k$) estimator $\tilde{f}_{k,p,b}$ for replicate $b$. Repeating this process independently for $b=1,\dots,B$ and aggregating via simple averaging yields the ensemble estimator
$\hat{f}_{k,m,p,B}=\frac{1}{B}\sum_{b=1}^{B}\tilde{f}_{k,p,b}.$
By the strong law of large numbers, the estimator $\hat{f}_{k,m,p,B}$ converges almost surely to its expected value as $B\to\infty$, with the expectation taken over the randomness of the subset choices. For the rest of this paper, we use and refer to this limit model as the EFS($k,m$) estimator, denoting it as $\hat{f}_{k,m,p}$.
2.3 Prediction error decomposition
In a fixed design setting, the complexity of an estimator $\hat{f}$ is often quantified via its degrees of freedom (hastie2009elements),
$\operatorname{df}(\hat{f})=\sum_{i=1}^{n}\frac{\operatorname{Cov}(y_{i},\hat{f}(\mathbf{x}_{i}))}{\sigma^{2}}.$ (2)
Letting $\mathbf{f}=\mathbf{X}\boldsymbol{\beta}$ and $\hat{\mathbf{f}}=\mathbf{X}\hat{\boldsymbol{\beta}}$ denote the true and fitted function values respectively, the expected mean squared prediction error of $\hat{f}$ can be decomposed into its expected training error and a constant multiple of its degrees of freedom,
[…]
We will therefore analyze how EFS($k,m$) compares to FS($k$) on each term on the right-hand side of (3).
3 Properties of EFS($k,m$) weights
3.1 EFS($k,m$) as reweighted OLS
For the rest of this paper, we assume that the features are orthogonal.
Assumption 3.1 (Orthogonality).
The design matrix satisfies $\mathbf{X}^{\top}\mathbf{X}/n=\mathbf{I}$.
This assumption is unlikely to describe most real datasets, but it allows us to derive exact asymptotic expressions for coefficient behavior and precisely characterize how the randomization from feature subsampling propagates through the FS($k$) algorithm. This enables direct comparisons with the broader regularization landscape, including ridge, lasso (tibshirani1996regression), elastic net (zou2005regularization), LARS (efron2004least), SCAD (fan2001variable), MCP (zhang2010nearly), and SureShrink (donoho1995adapting), all of which have been studied in the orthogonal setting.
Let $\hat{f}_{\operatorname{OLS}}=\sum_{j=1}^{p}\hat{\beta}_{j,p}x_{j,p}$ denote the ordinary least squares (OLS) estimator for $f$. Under Assumption 3.1, its coefficients can be expressed as inner products $\hat{\beta}_{j,p}=\langle\mathbf{y},\mathbf{x}_{\cdot j,p}\rangle$. It turns out that both the FS($k$) and EFS($k,m$) estimators can be written in terms of these OLS coefficients. For ease of notation, we establish an ordering convention for the features $x_{1,p},x_{2,p},\ldots,x_{p,p}$ based on these coefficients.
Definition 3.1 (Feature ordering).
The inner products $\hat{\beta}_{j,p}=\langle\mathbf{y},\mathbf{x}_{\cdot j,p}\rangle$ are reindexed and sorted by magnitude in descending order,
$|\hat{\beta}_{1,p}|\geq|\hat{\beta}_{2,p}|\geq\ldots\geq|\hat{\beta}_{p,p}|.$ (4)
Note that this ordering is random as it depends on the response noise. For simplicity, we assume that the inequalities in (4) are strict apart from a set of measure zero, which holds whenever the noise variables have a continuous density.
Using Definition 3.1 together with the equivalence of FS($k$) and BSS, we can write the FS($k$) estimator as
$\hat{f}_{k,p}=\sum_{j=1}^{p}\mathbf{1}(j\leq k)\hat{\beta}_{j,p}x_{j,p},$ (5)
which reflects the familiar fact that, in the orthogonal setting, FS($k$) (and BSS) selects the top-$k$ features most correlated with the response. The indicators $\mathbf{1}(j\leq k)$ can be thought of as weights that modulate the OLS solution. Similarly, the $k$-step ensemble estimator has the form
$\hat{f}_{k,m,p}=\sum_{j=1}^{p}w_{j}^{k,m,p}\hat{\beta}_{j,p}x_{j,p},$ (6)
where the weights $w_{1}^{k,m,p},w_{2}^{k,m,p},\ldots,w_{p}^{k,m,p}$ are determined solely by structural parameters ($j,k,m,p$) rather than the actual coefficient values $\hat{\beta}_{j,p}$. As $m,p\rightarrow\infty$ with a fixed scaling ratio, we will show that these weights are closely approximated by a logistic function with midpoint $k+1/2$ and growth rate $\log(1-m/p)$,
$w_{j}^{k,m,p}\approx\frac{1}{1+\big(1-\frac{m}{p}\big)^{k-j+1/2}}.$ (7)
3.2 Basic properties
Since $w_{j}^{k,m,p}=\mathbb{P}(x_{j,p}\in\mathcal{M}_{k})$, the weight sequence clearly takes values in $[0,1]$ and sums to $k$, the effective number of variables in the model. Unfortunately, the exact values of the weights involve complicated combinatorics and do not have simple closed-form expressions in general. Nonetheless, one can show that the sequence is non-increasing and, more importantly, derive a recurrence relation in terms of $j$, $k$, and $p$, which serves as our main technical tool for the subsequent results.
Lemma 3.1 (Monotonicity).
For all $j<j^{\prime}$ and $k,m,p\geq 1$, we have $w_{j}^{k,m,p}\geq w_{j^{\prime}}^{k,m,p}$.
Lemma 3.2 (Recurrence relation).
The weights satisfy the recurrence relation
$w_{j}^{k,m,p}=\frac{\binom{p-j}{m-1}}{\binom{p}{m}}+\frac{\binom{p-j}{m}}{\binom{p}{m}}w_{j}^{k-1,m,p-1}+\left(1-\frac{\binom{p-j+1}{m}}{\binom{p}{m}}\right)w_{j-1}^{k-1,m,p-1},$ (8)
where $w_{j}^{0,m,p}=w_{j}^{k,m,0}=w_{0}^{k,m,p}=0$.
Proof sketch of Lemma 3.1.
Define the events $\mathcal{E}_{1}=\{x_{j,p}\in\mathcal{M}_{k},\,x_{j-1,p}\notin\mathcal{M}_{k}\}$ and $\mathcal{E}_{2}=\{x_{j-1,p}\in\mathcal{M}_{k},\,x_{j,p}\notin\mathcal{M}_{k}\}$. It suffices to show $\mathbb{P}(\mathcal{E}_{1})\leq\mathbb{P}(\mathcal{E}_{2})$.
We construct a measure-preserving map $T$ from $\mathcal{E}_{1}$ to $\mathcal{E}_{2}$ as follows. Given a random seed $\boldsymbol{\mathcal{V}}\in\mathcal{E}_{1}$ where $x_{j,p}$ is selected at step $i$, define $T(\boldsymbol{\mathcal{V}})$ by swapping the roles of $x_{j,p}$ and $x_{j-1,p}$ in the candidate sets from step $i$ onward. Since the algorithm depends only on the ordinal structure of coefficients, this transformation preserves probabilities and maps $\mathcal{E}_{1}$ into $\mathcal{E}_{2}$, establishing the desired inequality. ∎
Proof sketch of Lemma 3.2.
Let $J$ denote the index of the feature selected at the first step. The key observation is that conditioning on $J$ reduces the problem to a smaller instance of EFS.
If $J=j$, then $x_{j,p}\in\mathcal{M}_{k}$ with certainty. If $J\neq j$, the remaining k1k-1 steps are equivalent to running EFS on p1p-1 features with reassigned indices. When $J>j$, the index $j$ remains unchanged, contributing $w_{j}^{k-1,m,p-1}$. When $J<j$, the index shifts to j1j-1, contributing $w_{j-1}^{k-1,m,p-1}$.
Computing $\mathbb{P}(J=j)$, $\mathbb{P}(J>j)$, and $\mathbb{P}(J<j)$ via combinatorics (specifically, the probability that $j$ is the minimum index in a uniformly random $m$-subset) and applying the hockey-stick identity to simplify the resulting sums yields the recurrence. ∎
3.3 Logistic approximation
The recurrence (8) is three-dimensional (in $j$, $k$, and $p$) and non-homogeneous, making it challenging to find a simple closed-form solution. However, in the scaling limit as $m\to\infty$, $p\to\infty$, and $m/p\to\gamma$ for some constant $\gamma\in[0,1]$, the recurrence converges to a two-dimensional system that further decouples into a one-dimensional subsystem with an explicit analytical solution.
Theorem 3.1 (Asymptotic weights for fixed $k$).
The limiting weights $\tilde{w}_{j}^{k,\gamma}=\lim\limits_{\begin{subarray}{c}m\to\infty,\,p\to\infty,\\ m/p\to\gamma\end{subarray}}w_{j}^{k,m,p}$
  1. (a)
    exist and equal
    $\sum_{i=1}^{k}(-1)^{k-i}\prod_{l=i}^{k}\big(e^{-\alpha(j-l)}-e^{-\alpha j}\big),$ (9)
  2. (b)
    satisfy the bounds
    $\frac{1}{1+e^{-\alpha\left(h(\alpha,k)-j\right)}}\leq\tilde{w}_{j}^{k,\gamma}\leq\frac{1}{1+e^{-\alpha\left(h(\alpha,k)+1-j\right)}},$ (10)
    where $h(\alpha,k)=\frac{1}{\alpha}\log(e^{\alpha k}-1)$ and $\alpha\in(0,\infty)$ is defined via $\gamma=1-e^{-\alpha}$.
Proof sketch.
The argument proceeds in four steps.
Step 1. We derive the limiting recurrence. Taking $m,p\to\infty$ with $m/p\to\gamma$ in the coefficients of (8) yields
$\frac{\binom{p-j}{m-1}}{\binom{p}{m}}\to(e^{\alpha}-1)e^{-\alpha j},\quad\frac{\binom{p-j}{m}}{\binom{p}{m}}\to e^{-\alpha j},\quad 1-\frac{\binom{p-j+1}{m}}{\binom{p}{m}}\to 1-e^{-\alpha(j-1)},$
where $\alpha=-\log(1-\gamma)$. This produces the limiting recurrence
$\tilde{w}_{j}^{k,\gamma}=(e^{\alpha}-1)e^{-\alpha j}+e^{-\alpha j}\tilde{w}_{j}^{k-1,\gamma}+(1-e^{-\alpha(j-1)})\tilde{w}_{j-1}^{k-1,\gamma}.$
Existence of the limits follows by induction on $k$, using $\tilde{w}_{j}^{0,\gamma}=0$ as the base case.
Step 2. We decouple the recurrence. Through algebraic manipulation, this two-dimensional recurrence separates into two independent one-dimensional recurrences,
$\tilde{w}_{j}^{k,\gamma}$ $=(e^{-\alpha(j-k)}-e^{-\alpha j})(1-\tilde{w}_{j}^{k-1,\gamma}),$
$\tilde{w}_{j}^{k,\gamma}$ $=1-e^{-\alpha k}-(e^{-\alpha(k-j+1)}-e^{-\alpha k})\tilde{w}_{j-1}^{k,\gamma}.$
Step 3. We solve for the closed form. The first recurrence is a first-order linear relation in $k$ with non-constant coefficients. Solving it directly yields the closed form (9).
Step 4. We establish the bounds. The weights satisfy $\tilde{w}_{j}^{k,\gamma}\leq\tilde{w}_{j}^{k+1,\gamma}$ (more steps can only increase inclusion probability). Combining this monotonicity with the first decoupled recurrence gives the lower bound. The upper bound follows by a symmetric argument using $\tilde{w}_{j}^{k,\gamma}\geq\tilde{w}_{j}^{k-1,\gamma}$. ∎
The bounds in (10) reveal that the weights are sandwiched between two logistic functions with midpoints $h(\alpha,k)$ and $h(\alpha,k)+1$ and common growth rate $\alpha=-\log(1-\gamma)$. When $\alpha k$ is large and $\gamma$ is small, we have $h(\alpha,k)\approx k$ and $\alpha\approx\gamma$, respectively.
These approximations clarify how the EFS($k,m$) parameters affect the shape of the weight function. (If sampling is performed with replacement, all asymptotic results hold with $\alpha$ replaced by $\gamma$.) As $k$ increases, the logistic curve translates rightward. As $\gamma$ increases (larger $m$), the curve becomes steeper, more closely approximating the sharp thresholding of FS($k$). Figure 1 illustrates this behavior for $\gamma\in\{0.05,0.15,0.333\}$.
Figure 1: Comparison of $\tilde{w}_{j}^{k,\gamma}$ (solid) and the logistic approximation (7) (dashed) for $\gamma=0.05,0.15,$ and $0.333$, with $k=10$.
The logistic approximation becomes even more precise, differing only by a half-index shift rather than a full-index shift, when the selection threshold $k$ also tends to infinity. We capture this by introducing the “offset” parameter d=kjd=k-j, measuring distance from the selection threshold.
Theorem 3.2 (Asymptotic weights for large $k$).
The limiting weights $\check{w}_{d}^{\gamma}=\lim\limits_{\begin{subarray}{c}k\to\infty\end{subarray}}\lim\limits_{\begin{subarray}{c}m\to\infty,\,p\to\infty,\\ m/p\to\gamma\end{subarray}}w_{k-d}^{k,m,p}$
  1. (a)
    exist and equal
    $\sum_{i=0}^{\infty}(-1)^{i}e^{-\alpha i(d+i/2+1/2)},$ (11)
  2. (b)
    satisfy the symmetry property $\check{w}_{d}^{\gamma}=1-\check{w}_{-(d+1)}^{\gamma}$,
  3. (c)
    satisfy the bounds
    1. (i)
      $\frac{1}{1+e^{-\alpha d}}\leq\check{w}_{d}^{\gamma}\leq\frac{1}{1+e^{-\alpha(d+1/2)}},\quad d<0$,
    2. (ii)
      $\frac{1}{1+e^{-\alpha(d+1/2)}}\leq\check{w}_{d}^{\gamma}\leq\frac{1}{1+e^{-\alpha(d+1)}},\quad d\geq 0$,
  4. (d)
    have a sigmoidal (S-shaped) pattern with respect to $d$.
Proof sketch.
We establish each part in turn.
Part (a). Using the recurrences from Theorem 3.1, one can show $\tilde{w}_{j+1}^{k,\gamma}-\tilde{w}_{j}^{k-1,\gamma}\geq 0$. For fixed offset d=kjd=k-j, the sequence $\{\tilde{w}_{k-d}^{k,\gamma}\}_{k\geq d+1}$ is therefore bounded and non-decreasing, hence convergent by the monotone convergence theorem. Taking $k\to\infty$ in the closed form (9) with j=kdj=k-d, careful analysis of convergence rates (tail terms decay exponentially) yields (11).
Part (b). Direct calculation using (11) verifies $1-\check{w}_{-(d+1)}^{\gamma}=\check{w}_{d}^{\gamma}$.
Part (c). These follow from taking $k\to\infty$ in the bounds (10).
Part (d). The bounds immediately imply that $\check{w}_{d}^{\gamma}$ is non-decreasing in $d$. Convexity for d<1/2d<-1/2 and concavity for d>1/2d>-1/2 are established by analyzing the second derivative of an extended continuous version of $\check{w}_{d}^{\gamma}$, using a $q$-series identity from coogan2003q. ∎
4 Main results
We now present our main results comparing EFS($k,m$) to FS($k$) in terms of the training error and degrees of freedom decomposition (3).
4.1 Expected training error
Let $\hat{m}$ denote the value of $m$ minimizing the training error $\|\mathbf{y}-\hat{\mathbf{f}}_{k,m,p}\|^{2}$. Since FS($k$) coincides with EFS($k,m$) when $m=p$, the expected training error of EFS($k,\hat{m}$) is trivially no larger than that of FS($k$). What is surprising is that $\hat{m}$ can take values less than $m$ and that we can quantify the gap in expected training error using the asymptotic formulas (10).
Theorem 4.1 (Training error improvement).
Grant Assumption 3.1 and fix $k\geq 1$. Assume $p$ is large enough that
$p\geq C\Bigg(\frac{\sum_{j=1}^{k}\mathbb{E}\big[\hat{\beta}^{2}_{j,p}\big]}{\sum_{j=k+1}^{2k}\mathbb{E}\big[\hat{\beta}^{2}_{j,p}\big]}\Bigg)^{3},$
where $C$ is a positive constant depending only on $k$. Then
$\mathbb{E}\big[\|\mathbf{y}-\hat{\mathbf{f}}_{k,\hat{m},p}\|^{2}\big]\leq\mathbb{E}\big[\|\mathbf{y}-\hat{\mathbf{f}}_{k,p}\|^{2}\big]-\frac{1}{4}\frac{\big(\frac{1}{k}\sum_{j=k+1}^{2k}\mathbb{E}\big[\hat{\beta}^{2}_{j,p}\big]\big)^{2}}{\frac{1}{k}\sum_{j=1}^{k}\mathbb{E}[\hat{\beta}^{2}_{j,p}]}.$ (12)
Proof sketch.
The argument proceeds in four steps.
Step 1. We decompose the training error gap. From (6) and (5), the gap decomposes as
$\|\mathbf{y}-\hat{\mathbf{f}}_{k,p}\|^{2}-\|\mathbf{y}-\hat{\mathbf{f}}_{k,m,p}\|^{2}=\underbrace{\sum_{j=k+1}^{p}\hat{\beta}_{j,p}^{2}\big(1-(1-w_{j}^{k,m,p})^{2}\big)}_{\text{gain from including }j>k}-\underbrace{\sum_{j=1}^{k}\hat{\beta}_{j,p}^{2}(1-w_{j}^{k,m,p})^{2}}_{\text{loss from excluding }j\leq k}.$
Step 2. We apply Chebyshev’s sum inequality. Both $\mathbb{E}[\hat{\beta}_{j,p}^{2}]$ and $w_{j}^{k,m,p}$ are non-increasing in $j$. Chebyshev’s inequality therefore bounds the loss term from above and the gain term from below by products of averages,
$\sum_{j=1}^{k}\mathbb{E}[\hat{\beta}_{j,p}^{2}](1-w_{j}^{k,m,p})^{2}$ $\leq\frac{1}{k}\sum_{j=1}^{k}\mathbb{E}[\hat{\beta}_{j,p}^{2}]\cdot\sum_{j=1}^{k}(1-w_{j}^{k,m,p})^{2},$
$\sum_{j=k+1}^{2k}\mathbb{E}[\hat{\beta}_{j,p}^{2}]\big(1-(1-w_{j}^{k,m,p})^{2}\big)$ $\geq\frac{1}{k}\sum_{j=k+1}^{2k}\mathbb{E}[\hat{\beta}_{j,p}^{2}]\cdot\sum_{j=k+1}^{2k}\big(1-(1-w_{j}^{k,m,p})^{2}\big).$
Step 3. We bound the weight sums using the logistic approximation. The bounds (10) imply $w_{j}^{k,\gamma}\leq 1/2$ for $j\geq k+1$, so $1-(1-w_{j}^{k,\gamma})^{2}\geq\frac{3}{2}w_{j}^{k,\gamma}$. Combined with the lower bound on $w_{j}^{k,\gamma}$, this gives $\sum_{j=k+1}^{2k}(1-(1-w_{j}^{k,\gamma})^{2})\gtrsim 1-\gamma$. A similar analysis yields $\sum_{j=1}^{k}(1-w_{j}^{k,\gamma})^{2}\lesssim(1-\gamma)^{2}$.
Step 4. We optimize over $\gamma$. Let $A=\frac{1}{k}\sum_{j=1}^{k}\mathbb{E}[\hat{\beta}_{j,p}^{2}]$ and $B=\frac{1}{k}\sum_{j=k+1}^{2k}\mathbb{E}[\hat{\beta}_{j,p}^{2}]$. The gap is at least $c_{1}B(1-\gamma)-c_{2}A(1-\gamma)^{2}$ for universal constants $c_{1},c_{2}>0$. Optimizing over $\gamma$ and verifying the condition on $p$ ensures the asymptotic approximations are valid, yields the bound $\frac{1}{4}B^{2}/A$. ∎
The bound (12) is larger when the coefficients $\hat{\beta}_{k+1,p},\hat{\beta}_{k+2,p},\ldots$ decay slowly. This makes intuitive sense, as in such settings spreading the weight budget more evenly across features improves the fit.
4.2 Degrees of freedom
We now show that, under a natural assumption, decreasing the feature subsampling ratio $m/p$ also decreases the degrees of freedom. The key is a decomposition of the degrees of freedom of EFS($k,m$) as a weighted sum over those of FS($j$) for varying $j$.
Theorem 4.2 (Degrees of freedom improvement).
Grant Assumption 3.1. For all $k,m,p\geq 1$, the degrees of freedom satisfy
$\operatorname{df}(\hat{f}_{k,m,p})=\sum_{j=1}^{p}w_{j}^{k,m,p}\left(\operatorname{df}(\hat{f}_{j,p})-\operatorname{df}(\hat{f}_{j-1,p})\right),$ (13)
where $\hat{f}_{0,p}=0$ and $\operatorname{df}(\hat{f}_{0,p})=0$. In particular, if $k\mapsto\operatorname{df}(\hat{f}_{k,p})$ is concave, then $m\mapsto\operatorname{df}(\hat{f}_{k,m,p})$ is increasing.
Proof sketch.
The decomposition (6) implies
$\hat{f}_{k,m,p}=\sum_{j=1}^{p}w_{j}^{k,m,p}(\hat{f}_{j,p}-\hat{f}_{j-1,p}).$
Since degrees of freedom is linear in the estimator, (13) follows immediately.
For the second statement, we establish that the weight sequences form a majorization hierarchy. For $m<m^{\prime}$,
$\sum_{j=1}^{l}w_{j}^{k,m,p}\leq\sum_{j=1}^{l}w_{j}^{k,m^{\prime},p}\quad\text{for all }l=1,\dots,p,$
with equality when $l=p$. The conclusion then follows from a standard majorization inequality (marshall11).
To prove the majorization, we construct a coupling between the EFS processes at levels $m$ and $m^{\prime}$. At each iteration, first draw a subset $B_{m}$ of size $m$, then augment it to $B_{m^{\prime}}\supseteq B_{m}$ of size $m^{\prime}$ by adding $m^{\prime}-m$ additional elements. Let $Z_{k,m,p}(l)=|S_{k,m,p}\cap[l]|$ count selected features with index at most $l$. Note that $\mathbb{E}[Z_{k,m,p}(l)]=\sum_{j=1}^{l}w_{j}^{k,m,p}$.
An inductive argument shows $Z_{k,m^{\prime},p}(l)\geq Z_{k,m,p}(l)$ for all $l$ and $k$. The key insight is that $\min B_{m^{\prime}}\leq\min B_{m}$, so the feature selected with parameter $m^{\prime}$ has index at most that selected with $m$. Careful case analysis completes the induction. ∎
The concavity condition in Theorem 4.2 is likely to hold in low-SNR scenarios where $|\hat{\beta}_{k,p}|\gg|\beta_{k,p}|$. Indeed, in such cases,
$\frac{\sigma^{2}}{n}\left(\operatorname{df}(\hat{f}_{k,p})-\operatorname{df}(\hat{f}_{k-1,p})\right)=\frac{\sigma^{2}}{n}\operatorname{df}\big(\hat{\beta}_{k,p}x_{k,p}\big)=\mathbb{E}\big[\hat{\beta}_{k,p}\big(\hat{\beta}_{k,p}-\beta_{k,p}\big)\big]\approx\mathbb{E}\big[\hat{\beta}_{k,p}^{2}\big],$ (14)
which is non-increasing in $k$ by Definition 3.1. Empirical support for the concavity of FS($k$) degrees of freedom, even in non-orthogonal settings, appears in Figures 1 and 4 of hastie2020extended and Figure 7 of mentch2020randomization.
Remark.
For a constant model ($\boldsymbol{\beta}=\mathbf{0}$), the approximation in (14) becomes exact.
When concavity fails, the degrees of freedom can be non-monotonic in $m$. This happens, for instance, when applying FS($k$) and EFS($k,m$) to data generated from an exactly $k$-sparse model.
Proposition 4.1 (Non-monotonic degrees of freedom).
Grant Assumption 3.1 and suppose $\boldsymbol{\beta}$ is exactly $k$-sparse with all nonzero coefficients equal to 1, and that the noise is Gaussian. Then
$\lim\limits_{\begin{subarray}{c}m,\,p,\,n\to\infty,\\ m/p\to\gamma\end{subarray}}\frac{\operatorname{df}(\hat{f}_{k,m,p})-\operatorname{df}(\hat{f}_{k,p})}{\log p}\geq C,$
where $C>0$ depends only on $\gamma$ and $k$.
Proof sketch.
Because of (13), we know that $\operatorname{df}(\hat{f}_{k,m,p})\geq w_{j}^{k,m,p}\operatorname{df}(\hat{\beta}_{k+1,p}x_{k+1,p})$. It hence suffices to show that $\operatorname{df}(\hat{\beta}_{k+1,p}x_{k+1,p})\approx\mathbb{E}[\hat{\beta}_{k+1,p}^{2}]$ is of order $\log p$. This follows from the fact that $\hat{\beta}_{k+1,p}x_{k+1,p}$ is essentially the maximum of pkp-k independent $\chi_{1}^{2}$ random variables. ∎
5 Further comparisons
5.1 Comparison with elastic net
It is instructive to compare EFS($k,m$) with elastic net regression (zou2005regularization), which combines the $\ell_{1}$ and $\ell_{2}$ penalties of lasso and ridge,
$\hat{\boldsymbol{\beta}}^{\text{EN}}=\operatorname*{arg\,min}_{\boldsymbol{\beta}\in\mathbb{R}^{p}}\Big\{\|\mathbf{y}-\mathbf{X}\boldsymbol{\beta}\|^{2}+\lambda_{1}\|\boldsymbol{\beta}\|_{\ell_{1}}+\lambda_{2}\|\boldsymbol{\beta}\|^{2}_{\ell_{2}}\Big\}.$
Under Assumption 3.1, elastic net takes a simple form (zou2005regularization),
$\hat{\beta}^{\text{EN}}_{j,p}=\frac{(|\hat{\beta}_{j,p}|-\lambda_{1}/2)_{+}}{1+\lambda_{2}}\operatorname{sgn}(\hat{\beta}_{j,p}).$ (15)
For comparison, Theorem 3.2 implies the EFS($k,m$) coefficient is approximately
$\hat{\beta}^{\text{EFS}}_{j,p}\approx\frac{\hat{\beta}_{j,p}}{1+\big(1-\frac{m}{p}\big)^{k-j+1/2}}.$
Elastic net was originally designed to mitigate lasso’s tendency toward highly sparse solutions, which causes instability when features are correlated. From (15), we see that even in the orthogonal setting, elastic net spreads the parameter budget across more coefficients by shrinking all nonzero coefficients by the constant factor $1/(1+\lambda_{2})$. This is similar in spirit to feature subsampling in EFS($k,m$), but the families of attainable solutions differ, yielding different predictive performances on any given dataset.
5.2 Revisiting best subset vs. lasso
hastie2020extended compared BSS and lasso, showing that BSS (and FS($k$)) achieves higher accuracy than lasso in high-SNR settings, and that lasso dominates in low-SNR regimes. They explain this by observing that “different procedures bring us from the high-bias to the high-variance ends of the tradeoff along different model paths,” with neither algorithm’s solution path completely dominating the other. See also ghosh2025signal for recent theoretical support.
Our results add to this discussion by showing that the BSS solution path is not Pareto optimal. EFS can simultaneously reduce both bias and variance. The ensemble weights trace yet another solution path, distinct from those of lasso and elastic net.
5.3 Shrinkage
Decreasing $m$ does not shrink every coefficient (some weights increase and others decrease), but it does shrink the fitted values toward zero.
Proposition 5.1.
Grant Assumption 3.1. For $m<m^{\prime}$ and any $k,p\geq 1$,
$\|\hat{\mathbf{f}}_{k,m,p}\|\leq\|\hat{\mathbf{f}}_{k,m^{\prime},p}\|.$
Proof sketch.
From (6), it suffices to show $\sum_{j=1}^{p}\hat{\beta}_{j,p}^{2}(w_{j}^{k,m,p})^{2}\leq\sum_{j=1}^{p}\hat{\beta}_{j,p}^{2}(w_{j}^{k,m^{\prime},p})^{2}$. This follows from the majorization established in Theorem 4.2 together with the Schur-convexity of $g(w_{1},\ldots,w_{p})=\sum_{j}\beta_{j}^{2}w_{j}^{2}$. ∎
The extremes are as follows. Maximal shrinkage occurs at $m=1$, where $\|\hat{\mathbf{f}}_{k,1,p}\|=(k/p)\|\hat{\boldsymbol{\beta}}\|_{\ell_{2}}$. No shrinkage occurs at $m=p$, where EFS($k,m$) reduces to FS($k$).
6 Another mechanism for training error reduction
Theorem 4.1 establishes training error reduction via the “enlarged model class” mechanism of curth2024random. The orthogonality assumption makes FS($k$) equivalent to BSS, so randomization helps only by searching over a richer family of coefficient profiles. However, liu2025randomization identified a second mechanism. Feature subsampling can help individual base estimators escape local optima that greedy optimization would otherwise get trapped in.
We now construct an example demonstrating this second mechanism. In the orthogonal setting, FS($k$) always selects the $k$ features most correlated with the response, which means that individual EFS base estimators cannot outperform it. We therefore introduce correlation in the form of a “spurious” feature $x_{p,p}$ that is strongly correlated with the response but not part of the optimal representation. Greedy FS selects this spurious feature, but EFS base estimators may avoid it if $x_{p,p}$ is excluded from the first candidate set.
Proposition 6.1 (Escaping local minima).
Consider a design matrix $\mathbf{X}$ with $\langle\mathbf{x}_{\cdot i,p},\mathbf{x}_{\cdot j,p}\rangle=\delta_{ij}$ for $1\leq i,j\leq{p-1}$, $\langle\mathbf{x}_{\cdot i,p},\mathbf{x}_{\cdot p,p}\rangle=3^{-1/2}$ for $i\in\{1,2\}$, and $\langle\mathbf{x}_{\cdot i,p},\mathbf{x}_{\cdot p,p}\rangle=0$ otherwise. Assume the noiseless model $y=\beta(x_{1,p}+x_{2,p})+\zeta\sum_{i=3}^{p-1}x_{i,p}$ with $6^{-1/2}\beta>\zeta>0$. Let $\check{f}_{k,m,p}$ denote a single EFS base estimator. There exists $C>0$ depending only on $\beta,\xi$ such that for $k,p\geq C$,
$\mathbb{E}\big[\|\mathbf{y}-\hat{\mathbf{f}}_{k,p}\|^{2}\big]>\mathbb{E}\big[\|\mathbf{y}-\check{\mathbf{f}}_{k,m,p}\|^{2}\big].$
7 Simulations
To validate Theorems 4.1 and 4.2 and assess their generality beyond orthogonal designs, we conducted a simulation study. We generated IID training data $\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{n}$ with $n=1000$, $p=100$, where $\mathbf{x}_{i}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{\Sigma})$ with banded correlation $\Sigma_{ij}=\rho^{|i-j|}$ for $\rho=0.5$, $y_{i}=\boldsymbol{\beta}^{\top}\mathbf{x}_{i}+\varepsilon_{i}$ with $\varepsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$, and exact sparsity $\beta_{j}=\mathbf{1}(j\leq 10)$. The noise variance $\sigma^{2}$ was chosen to achieve the desired signal-to-noise ratio $\textnormal{SNR}=\boldsymbol{\beta}^{\top}\boldsymbol{\Sigma}\boldsymbol{\beta}/\sigma^{2}$.
For EFS, we used $B=500$ base learners and varied $m$ over a geometric grid from $2$ to $p$, selecting the optimal $m$ via 10-fold cross-validation for each $k\in[20]$. We compared degrees of freedom and training error for EFS($k,m$) against FS($k$) at each model size $k$.
Figure 2 shows results for $\textnormal{SNR}\in\{0.053,0.25,1.0\}$. At low SNR ($0.053$ and $0.25$), EFS($k,m$) reduces both degrees of freedom and training error relative to FS($k$) at each $k$, shifting the entire bias-variance tradeoff curve toward the origin. At high SNR ($1.0$), EFS($k,m$) performs worse near the true sparsity level $k=10$, consistent with Proposition 4.1.
(a) SNR =0.053=0.053
(b) SNR =0.25=0.25
(c) SNR =1.0=1.0
Figure 2: Degrees of freedom vs. training MSE for EFS($k,m$) and FS($k$) with $p/n=0.1$ and banded exact sparsity, as $k$ varies over $\{1,2,\ldots,20\}$ (true sparsity is 10). At each $k$, EFS($k,m$) reduces both quantities, shifting the tradeoff curve toward the origin. Gains are more pronounced at low SNR. At high SNR, EFS($k,m$) and FS($k$) become similar.
8 Discussion
We have revisited the role of feature subsampling in randomized ensembles such as random forests, focusing on subtleties that emerge when subsampling is interleaved with greedy optimization rather than applied to base learners fit via convex optimization. By analyzing ensemble forward selection under orthogonal fixed design, we moved beyond the empirical observations of curth2024random; liu2025randomization; revelas2025random to prove that EFS($k,m$) can simultaneously achieve smaller bias and smaller variance than the base estimator FS($k$). Our work is similar in spirit to mei2024exogenous, who analyzed an oracle version of RFs, but differs in our ability to analyze a non-oracle algorithm, explicitly quantify the performance gap, and characterize the resulting solution. In particular, we reveal a logistic reweighting of the OLS coefficients.
These results, alongside recent related work, broaden our understanding of randomization in RFs. mentch2020randomization concluded that feature subsampling is mainly useful in low-SNR regimes, but our results show it can also help in the absence of noise (see Proposition 6.1). Moreover, since feature subsampling is often viewed as implicit regularization, our findings urge closer attention to the nuances of this phenomenon. In particular, mental models derived from ridge and lasso regression may not transfer cleanly to algorithmic regularization via feature subsampling, dropout, or gradient descent (wager2013dropout; neyshabur2015searchrealinductivebias; arora2019implicit; bartlett2020benign).
The main theoretical limitation of our work is the orthogonality assumption. Our simulations partially bridge this gap, showing that EFS continues to improve upon FS in non-orthogonal settings. Extending the theory to handle correlated features is a promising direction for future work. For instance, one could study how structured perturbations of $\mathbf{X}^{\top}\mathbf{X}/n$ affect the ensemble weights. Additionally, our analysis applies directly only to forward selection. Extending the theoretical framework of mei2024exogenous to characterize how feature subsampling improves actual RFs remains an important open problem.
Reproducibility
The code used to reproduce all numerical experiments in this paper is available at
 github.com/yuyumi/EFS.
Acknowledgements
J.M.K. was partially supported from the National Science Foundation (NSF) through CAREER DMS-2239448. Y.S.T. was supported from NUS Startup Grant A-8000448-00-00 and MOE AcRF Tier 1 Grant A-8002498-00-00.
Supplement to “Revisiting Randomization in Greedy Model Search”
S.1 Proofs
S.1.1 Proofs for finite EFS($k,m$) weights
Proof of Lemma 3.1.
To prove monotonicty of the weights with respect to $j$, recall that $w_{j}^{k,m,p}=\mathbb{P}(x_{j,p}\in\mathcal{M}_{k})$. Here, the randomness is over the choices of the sets, $\boldsymbol{{\mathcal{V}}}=\{\mathcal{V}_{1},\mathcal{V}_{2},\ldots,\mathcal{V}_{k}\}$. Define the events
$\mathcal{E}_{1}=\{x_{j,p}\in\mathcal{M}_{k},\;x_{j-1,p}\notin\mathcal{M}_{k}\},$
and
$\mathcal{E}_{2}=\{x_{j-1,p}\in\mathcal{M}_{k},\;x_{j,p}\notin\mathcal{M}_{k}\}.$
It suffices to show that $\mathbb{P}(\mathcal{E}_{1})\leq\mathbb{P}(\mathcal{E}_{2})$. To show this, we will define a measure-preserving transformation $T\colon\mathcal{E}_{1}\to\mathcal{E}_{2}$ with $T(\mathcal{E}_{1})\subset\mathcal{E}_{2}$. Assuming such a transformation exists, we have
$\mathbb{P}(\mathcal{E}_{2})\geq\mathbb{P}(T(\mathcal{E}_{1}))=\mathbb{P}(\mathcal{E}_{1}).$
[…]
Proof of Lemma 3.2.
Let $J$ be denote the index of the feature selected at the first step. To derive the recurrence for $w_{j}^{k,m,p}$, we condition on the first selected element $x_{i,p}$, i.e., condition on the event $J=i$.
  1. 1.
    If $i=j$, which occurs with probability $\mathbb{P}(J=j)$, the element $x_{j,p}$ is selected in the first step. As a result, $x_{j,p}$ will always be included in $\mathcal{M}_{k}$.
  2. 2.
    If $i\neq j$, the remaining k1k-1 steps are equivalent to running EFS on p1p-1 features. The coefficients are reassigned as follows:
    $\hat{\beta}_{l,p-1}$ $\leftarrow\hat{\beta}_{l,p}\qquad\text{for }l<i$
    $\hat{\beta}_{l,p-1}$ $\leftarrow\hat{\beta}_{l+1,p}\quad\text{for }l\geq i,$
    where $l=1,2,\dots,p-1$. We then consider two subcases:
    1. (a)
      If $i>j$, which occurs with probability $\mathbb{P}(J>j)$, the index $j$ remains unchanged in the relative ranking. The weight contribution becomes $w_{j}^{k-1,m,p-1}$.
    2. (b)
      If $i<j$, which occurs with probability $\mathbb{P}(J<j)$, the index $j$ is shifted to the (j1)(j-1)-th position in the new ordering. The weight contribution becomes $w_{j-1}^{k-1,m,p-1}$.
At the first step of EFS, we select the feature with the largest absolute inner product $|\hat{\beta}_{j,p}|=|\langle\mathbf{y},\mathbf{x}_{\cdot j,p}\rangle|$ from an $m$-sized subset of $p$ features. Because there are no ties in the feature rankings, this is equivalent to selecting the minimum index $j$ in this subset. The probability of $x_{j,p}$ being selected is thus the probability that $j$ is the minimum index in a random $m$-sized subset of $p$ elements, which is $\binom{p-j}{m-1}/\binom{p}{m}$. Therefore, by the law of total probability,
$w_{j}^{k,m,p}$ $=\mathbb{P}(J=j)+w_{j}^{k-1,m,p-1}\mathbb{P}(J>j)+w_{j-1}^{k-1,m,p-1}\mathbb{P}(J<j)$ (S.1)
$=\frac{\binom{p-j}{m-1}}{\binom{p}{m}}+w_{j}^{k-1,m,p-1}\sum^{p}_{i=j+1}\frac{\binom{p-i}{m-1}}{\binom{p}{m}}+w_{j-1}^{k-1,m,p-1}\sum^{j-1}_{i=1}\frac{\binom{p-i}{m-1}}{\binom{p}{m}}$.
Now, by the hockey-stick identity from combinatorics, we can simplify the coefficients from (S.1):
$\sum^{p}_{i=j+1}\binom{p-i}{m-1}$ $=\binom{p-j}{m},\text{ and}$
$\sum^{j-1}_{i=1}\binom{p-i}{m-1}$ $=\binom{p}{m}-\binom{p-j+1}{m}$.
Plugging in these two expressions into (S.1) gives us recurrence (8). ∎
S.1.2 Proofs for asymptotic EFS($k,m$) weights
To prove Theorem 3.1, we give several useful lemmas. The first lemma shows that for fixed $j$ and $k$, we have the following existence and recurrence.
Lemma S.1.1.
The limit $\tilde{w}_{j}^{k,\gamma}=\lim\limits_{\begin{subarray}{c}m\to\infty,\,p\to\infty,\\ m/p\to\gamma\end{subarray}}w_{j}^{k,m,p}$ exists and satisfies the following recurrence for $j\geq 1$,
$\tilde{w}_{j}^{k,\gamma}=(e^{\alpha}-1)e^{-\alpha j}+e^{-\alpha j}\tilde{w}_{j}^{k-1,\gamma}+\big(1-e^{-\alpha(j-1)}\big)\tilde{w}_{j-1}^{k-1,\gamma}.$ (S.2)
However, recurrence (S.2) contains dependencies on both $\tilde{w}_{j}^{k-1,\gamma}$ and $\tilde{w}_{j-1}^{k-1,\gamma}$. These dependencies make it difficult to construct a closed form and uncover properties of $\tilde{w}_{j}^{k,\gamma}$. Ideally, we would like recurrences involving a change in $j$ to be separate from recurrences involving a change in $k$. Fortunately, recurrence (S.2) can be separated into the following pair of recurrences.
Lemma S.1.2.
From the the recurrence in Lemma S.1.1, we can construct the following recurrence pair for $j\geq 1$:
$\tilde{w}_{j}^{k,\gamma}=\big(e^{-\alpha(j-k)}-e^{-\alpha j}\big)\big(1-\tilde{w}_{j}^{k-1,\gamma}\big),\quad\tilde{w}_{j}^{0,\gamma}=0,$ (S.3)
$\tilde{w}_{j}^{k,\gamma}=1-e^{-\alpha k}-\big(e^{-\alpha(k-j+1)}-e^{-\alpha k}\big)\tilde{w}_{j-1}^{k,\gamma},\quad\tilde{w}_{0}^{k,\gamma}=0.$ (S.4)
Proof of Lemma S.1.1.
To show the limit form of the coefficients in recurrence (S.2), we derive the following result using basic asymptotic facts. For a fixed $j$, for $m,p\to\infty$ and $m/p\to\gamma$, we have
$\frac{\binom{p-j}{m-1}}{\binom{p}{m}}=\frac{m}{p-j-m+1}\frac{\binom{p-j}{m}}{\binom{p}{m}}=\frac{m}{p-j-m+1}\prod^{j}_{i=1}\frac{i+p-m}{i+p}\underset{m,p\to\infty}{\longrightarrow}\frac{\gamma}{1-\gamma}(1-\gamma)^{j}=(e^{\alpha}-1)e^{-\alpha j}.$
Applying the same method to the other coefficients in the recurrence gives us the following:
$\frac{\binom{p-j}{m}}{\binom{p}{m}}$ $\underset{m,p\to\infty}{\longrightarrow}e^{-\alpha j},\text{ and}$
$\left(1-\frac{\binom{p-j+1}{m}}{\binom{p}{m}}\right)$ $\underset{m,p\to\infty}{\longrightarrow}\big(1-e^{-\alpha(j-1)}\big).$
Now, we show that $\tilde{w}_{j}^{k,\gamma}$ is well defined via induction on $k$. For the base case, let $k=1$. Note that $w_{j}^{0,m,p}=0$ for all $j$ and all $p$, so
$\tilde{w}_{j}^{0,\gamma}=\lim\limits_{\begin{subarray}{c}m\to\infty,\,p\to\infty,\\ m/p\to\gamma\end{subarray}}w_{j}^{0,m,p}=0.$
Therefore, $\tilde{w}_{j}^{0,\gamma}$ is well defined for all $j$. Now suppose that $\tilde{w}_{j}^{k,\gamma}$ is well defined for all steps up to $k$ and for all $j$. Consider step $(k+1)$. From recurrence (8), we have
$\lim\limits_{\begin{subarray}{c}m\to\infty,\,p\to\infty,\\ m/p\to\gamma\end{subarray}}w_{j}^{k+1,m,p}$
$=\lim\limits_{\begin{subarray}{c}m\to\infty,\,p\to\infty,\\ m/p\to\gamma\end{subarray}}\left(\frac{\binom{p-j}{m-1}}{\binom{p}{m}}+\frac{\binom{p-j}{m}}{\binom{p}{m}}w_{j}^{k,m,p}+w_{j-1}^{k,m,p}\left(1-\frac{\binom{p-j+1}{m}}{\binom{p}{m}}\right)\right)$
$=(e^{\alpha}-1)e^{-\alpha j}+e^{-\alpha j}\Bigg(\lim\limits_{\begin{subarray}{c}m\to\infty,\,p\to\infty,\\ m/p\to\gamma\end{subarray}}w_{j}^{k,m,p}\Bigg)+\big(1-e^{-\alpha(j-1)}\big)\Bigg(\lim\limits_{\begin{subarray}{c}m\to\infty,\,p\to\infty,\\ m/p\to\gamma\end{subarray}}w_{j-1}^{k,m,p}\Bigg)$
$=(e^{\alpha}-1)e^{-\alpha j}+e^{-\alpha j}\tilde{w}_{j}^{k,\gamma}+\big(1-e^{-\alpha(j-1)}\big)\tilde{w}_{j-1}^{k,\gamma}.$
Since $\tilde{w}_{j}^{k,\gamma}$ and $\tilde{w}_{j-1}^{k,\gamma}$ exist for all $j$ by inductive assumption, $w_{j}^{k+1,\gamma}$ exists as well for all $j$. Therefore, by induction, we know that the limit $\tilde{w}_{j}^{k,\gamma}$ exists for all $j,k$ and satisfies the recurrence in Lemma S.1.1. ∎
Proof of Lemma S.1.2.
We will induct on $k$. First, we verify that both recurrences satisfy the base case of $k=1$. Note that from previous arguments,
$\tilde{w}_{j}^{1,\gamma}$ $=\lim\limits_{\begin{subarray}{c}m\to\infty,\,p\to\infty,\\ m/p\to\gamma\end{subarray}}\frac{\binom{p-j}{m-1}}{\binom{p}{m}}=(e^{\alpha}-1)e^{-\alpha j}.$
For recurrence (S.3), we plug in $\tilde{w}_{j}^{0,\gamma}=0$ on the right hand side:
$\tilde{w}_{j}^{1,\gamma}$ $=\big(e^{-\alpha(j-1)}-e^{-\alpha j}\big)\big(1-\tilde{w}_{j}^{0,\gamma}\big)=(e^{\alpha}-1)e^{-\alpha j}.$
Thus, the recurrence (S.3) satisfies the base case $k=1$. Now, note that $\tilde{w}_{j-1}^{1,\gamma}=(e^{\alpha}-1)e^{-\alpha(j-1)}$. Plugging in this expression into recurrence (S.4), we see that
$\tilde{w}_{j}^{1,\gamma}=1-e^{-\alpha}-\big(e^{-\alpha(1-j+1)}-e^{-\alpha}\big)\tilde{w}_{j-1}^{1,\gamma}=(e^{\alpha}-1)e^{-\alpha j}.$
Thus, both recurrences satisfy the base case.
Now, for a fixed $j$, suppose that both recurrences are valid for steps up to $k$.
Consider step $(k+1)$. We show that $\tilde{w}_{j}^{k+1,\gamma}$ satisfies recurrence (S.3) in the following way. We start with the recurrence from Lemma S.1.1:
$\tilde{w}_{j}^{k+1,\gamma}$ $=(e^{\alpha}-1)e^{-\alpha j}+e^{-\alpha j}\tilde{w}_{j}^{k,\gamma}+\big(1-e^{-\alpha(j-1)}\big)\tilde{w}_{j-1}^{k,\gamma}.$
Substituting $\tilde{w}_{j-1}^{k,\gamma}$ using recurrence (S.4) under the inductive assumption, we obtain the following result:
$\tilde{w}_{j}^{k+1,\gamma}$ $=(e^{\alpha}-1)e^{-\alpha j}+e^{-\alpha j}\tilde{w}_{j}^{k,\gamma}+\big(1-e^{-\alpha(j-1)}\big)\left(\frac{1-e^{-\alpha k}-\tilde{w}_{j}^{k,\gamma}}{e^{-\alpha(k-j+1)}-e^{-\alpha k}}\right)$
$=\big(e^{-\alpha(j-k-1)}-e^{-\alpha j}\big)\big(1-\tilde{w}_{j}^{k,\gamma}\big).$
Therefore, recurrence (S.3) is satisfied for step $(k+1)$. Now, we prove that recurrence (S.4) is satisfied for step $(k+1)$ in a similar fashion. We plug in recurrence (S.3) into Lemma S.1.1:
$\tilde{w}_{j}^{k+1,\gamma}$ $=(e^{\alpha}-1)e^{-\alpha j}+e^{-\alpha j}\left(1-\frac{\tilde{w}_{j}^{k+1,\gamma}}{e^{-\alpha(j-k-1)}-e^{-\alpha j}}\right)$
$\qquad+\big(1-e^{-\alpha(j-1)}\big)\left(\frac{e^{-\alpha(j-k-2)}-e^{-\alpha(j-1)}-\tilde{w}_{j-1}^{k+1,\gamma}}{e^{-\alpha(j-k-2)}-e^{-\alpha(j-1)}}\right).$
Rearranging terms, we obtain the following recurrence:
$\tilde{w}_{j}^{k+1,\gamma}$ $=1-e^{-\alpha(k+1)}-\big(e^{-\alpha(k-j+2)}-e^{-\alpha(k+1)}\big)\tilde{w}_{j-1}^{k+1,\gamma}.$
Thus, recurrence (S.4) is satisfied for step $(k+1)$. Therefore, by induction, for a fixed $j$, both recurrences are satisfied for all $k\in\mathbb{N}^{+}$. ∎
Next, we prove Theorem 3.1.
Proof of Theorem 3.1.
To prove property (a), note that from Lemma S.1.1, we have that $\tilde{w}_{j}^{k,\gamma}$ exists. From (S.3), a first-order non-homogeneous recurrence relation with feature coefficients, we directly obtain the following closed form for $\tilde{w}_{j}^{k,\gamma}$ (elaydi2005introduction, Equation 1.2.4):
$\tilde{w}_{j}^{k,\gamma}=\sum_{i=1}^{k}(-1)^{k-i}\prod_{l=i}^{k}\big(e^{-\alpha(j-l)}-e^{-\alpha j}\big).$
To prove property (b), we first show the lower bound. According to Lemma S.1.5, we have $\tilde{w}_{j}^{k,\gamma}\geq\tilde{w}_{j}^{k-1,\gamma}$. From (S.3), we know
$\tilde{w}_{j}^{k,\gamma}$ $=\big(e^{-\alpha(j-k)}-e^{-\alpha j}\big)\big(1-\tilde{w}_{j}^{k-1,\gamma}\big)$
$\geq\big(e^{-\alpha(j-k)}-e^{-\alpha j}\big)\big(1-\tilde{w}_{j}^{k,\gamma}\big).$
Therefore, solving for $\tilde{w}_{j}^{k,\gamma}$, we have
$\tilde{w}_{j}^{k,\gamma}$ $\geq\frac{1-e^{-\alpha k}}{1-e^{-\alpha k}+e^{-\alpha(k-j)}}.$
Alternatively, we can rewrite recurrence (S.3) as
$\big(e^{-\alpha(j-k-1)}-e^{-\alpha j}\big)\big(1-\tilde{w}_{j}^{k,\gamma}\big)=\tilde{w}_{j}^{k+1,\gamma}\geq\tilde{w}_{j}^{k,\gamma}.$
Solving for $\tilde{w}_{j}^{k,\gamma}$, we have
$\tilde{w}_{j}^{k,\gamma}$ $\leq\frac{1-e^{-\alpha k}}{1-e^{-\alpha k}+e^{-\alpha(k-j+1)}}.$
Finally, note that we can write
$\begin{split}\frac{1-e^{-\alpha k}}{1-e^{-\alpha k}+e^{-\alpha(k-j)}}&=\frac{1}{1+\left(1-e^{-\alpha k}\right)^{-1}e^{-\alpha(k-j)}}\\ &=\frac{1}{1+\left(e^{\alpha k}-1\right)^{-1}e^{aj}}\\ &=\frac{1}{1+e^{-\alpha(h(a,k)-j)}}.\qed\end{split}$ (S.5)
Remark.
We give a remark showing that the weights $\tilde{w}_{j}^{k,\gamma}$ satisfy the following two additional properties. The weights $\tilde{w}_{j}^{k,\gamma}$
  1. (i)
    are non-increasing in $j$ for any given values of $k$ and $\gamma$, i.e., $\tilde{w}_{j}^{k,\gamma}\geq\tilde{w}_{j+1}^{k,\gamma}$,
  2. (ii)
    have sum $\sum_{j=1}^{\infty}\tilde{w}_{j}^{k,\gamma}=k$.
Property (i) directly follows from Lemma 3.2 by taking $m,p\to\infty$ and $m/p\to\gamma$.
To prove property (ii), note that from (S.2) we have
$\sum_{j=1}^{\infty}\tilde{w}_{j}^{k,\gamma}$ $=\sum_{j=1}^{\infty}(e^{\alpha}-1)e^{-\alpha j}+\sum_{j=1}^{\infty}\tilde{w}_{j}^{k-1,\gamma}+\sum_{j=1}^{\infty}e^{-\alpha j}\tilde{w}_{j}^{k-1,\gamma}-\sum_{j=1}^{\infty}e^{-\alpha(j-1)}\tilde{w}_{j-1}^{k-1,\gamma}$
$=1+\sum_{j=1}^{\infty}\tilde{w}_{j}^{k-1,\gamma}$
$=k,$
where we use the fact that $\tilde{w}_{j}^{0,\gamma}=\tilde{w}_{0}^{k,\gamma}=0$.
We now turn to the proof of Theorem 3.2. Before that, we present a useful lemma regarding an alternative form of $\tilde{w}_{j}^{k,\gamma}$.
Lemma S.1.3.
The asymptotic weights $\tilde{w}_{j}^{k,\gamma}$ can be rewritten in the following closed form
$\tilde{w}_{j}^{k,\gamma}$ $=\sum^{k}_{i=1}(-1)^{i+1}e^{\alpha i((k-j)-i/2+1/2)}\prod^{k}_{l=k-i+1}\big(1-e^{-l\alpha}\big).$
Proof of Lemma S.1.3.
Using the closed form from (9), we re-index the sum to obtain
$\tilde{w}_{j}^{k,\gamma}$ $=\sum_{i=1}^{k}(-1)^{k-i}\prod_{l=i}^{k}\big(e^{-\alpha(j-l)}-e^{-\alpha j}\big)$
$=\sum_{i=0}^{k-1}(-1)^{k-i+1}e^{-(k-i)j\alpha}\prod_{l=i+1}^{k}\big(e^{l\alpha}-1\big).$
Now, we note that
$\prod^{k}_{l=i+1}(e^{l\alpha}-1)$ $=e^{\alpha\left(\frac{k(k+1)}{2}-\frac{i(i+1)}{2}\right)}\prod^{k}_{l=i+1}\big(1-e^{-l\alpha}\big).$
Substituting this expression back into $\tilde{w}_{j}^{k,\gamma}$, we have
$\tilde{w}_{j}^{k,\gamma}$ $=\sum^{k-1}_{i=0}(-1)^{k-i+1}e^{-(k-i)j\alpha+\alpha\left(\frac{k(k+1)}{2}-\frac{i(i+1)}{2}\right)}\prod^{k}_{l=i+1}\big(1-e^{-l\alpha}\big)$
$=\sum^{k-1}_{i=0}(-1)^{k-i+1}e^{\alpha(k-i)\left(k-j-\frac{k-i}{2}+\frac{1}{2}\right)}\prod^{k}_{l=i+1}\big(1-e^{-l\alpha}\big).$
Finally, we re-index by setting $i$ to be kik-i, yielding
$\tilde{w}_{j}^{k,\gamma}$ $=\sum^{k}_{i=1}(-1)^{i+1}e^{\alpha i((k-j)-i/2+1/2)}\prod^{k}_{l=k-i+1}\big(1-e^{-l\alpha}\big).\qed$
We start to prove Theorem 3.2. For the following proofs, we perform the substitution d=kjd=k-j.
Proof of Theorem 3.2.
To prove property (a), we first prove the existence and derive a closed form for $\check{w}_{d}^{\gamma}$ slightly different from (11).
Lemma S.1.4.
The limit $\check{w}_{d}^{\gamma}=\lim\limits_{\begin{subarray}{c}k\to\infty,\,j\to\infty,\\ k-j=d\end{subarray}}\tilde{w}_{j}^{k,\gamma}$ exists and has the following closed form
$\check{w}_{d}^{\gamma}=\sum^{\infty}_{i=1}(-1)^{i+1}e^{\alpha i(d-i/2+1/2)}.$ (S.6)
Proof of Lemma S.1.4.
We begin by demonstrating the existence of the limit. To establish this, we first present a simple lemma regarding the monotonicity relation.
Lemma S.1.5.
For any given values of $j$ and $\gamma$, we have $\tilde{w}_{j}^{k,\gamma}\leq\tilde{w}_{j}^{k+1,\gamma}$ for all $k\geq 0$.
Proof of Lemma S.1.5.
The proof follows directly from the following inequality:
$\tilde{w}_{j}^{k,\gamma}=\lim\limits_{\begin{subarray}{c}m\to\infty,\,p\to\infty,\\ m/p\to\gamma\end{subarray}}\mathbb{P}(x_{j,p}\in\mathcal{M}_{k})\leq\lim\limits_{\begin{subarray}{c}m\to\infty,\,p\to\infty,\\ m/p\to\gamma\end{subarray}}\mathbb{P}(x_{j,p}\in\mathcal{M}_{k+1})=w_{j}(k+1,\gamma).\qed$
Note that by substituting (S.3) into (S.4), we have
$\tilde{w}_{j+1}^{k,\gamma}$ $=1-e^{-\alpha k}-e^{-\alpha(k-j)}\big(e^{-\alpha(j-k)}-e^{-\alpha j}\big)\big(1-\tilde{w}_{j}^{k-1,\gamma}\big)+e^{-\alpha k}\tilde{w}_{j}^{k,\gamma}$
$=e^{-\alpha k}w_{j}(k,\gamma)+\big(1-e^{-\alpha k}\big)\tilde{w}_{j}^{k-1,\gamma}.$
Thus, $\tilde{w}_{j+1}^{k,\gamma}-\tilde{w}_{j}^{k-1,\gamma}=e^{-\alpha k}\big(\tilde{w}_{j}^{k,\gamma}-\tilde{w}_{j}^{k-1,\gamma}\big)\geq 0$ for all $j,k\geq 1$, where we use Lemma S.1.5. Fix $d$, let j=kdj=k-d, and consider the sequence $\big\{\tilde{w}_{k-d}^{k,\gamma}\big\}^{\infty}_{k=d+1}$. This sequence is bounded and non-decreasing; thus, by the monotone convergence theorem, the limit of $\big\{\tilde{w}_{k-d}^{k,\gamma}\big\}^{\infty}_{k=d+1}$ exists. Hence, $\check{w}_{d}^{\gamma}$ exists and we must now identify its limit.
We will prove the closed form (S.6) by showing that for every $\epsilon>0$, there exists some $K$ such that for all $k>K$,
$\bigg|\tilde{w}_{k-d}^{k,\gamma}-\sum^{k}_{i=1}(-1)^{i+1}e^{\alpha i(d-i/2+1/2)}\bigg|<\epsilon.$
First, fix $\epsilon$ and suppose $k$ is even. We begin by using the triangle inequality, substituting in the closed form for $\tilde{w}_{k-d}^{k,\gamma}$ we obtained in Lemma S.1.3:
$\bigg|\tilde{w}_{k-d}^{k,\gamma}-\sum^{k}_{i=1}(-1)^{i+1}e^{\alpha i(d-i/2+1/2)}\bigg|$
$=\bigg|\sum^{k}_{i=1}(-1)^{i+1}e^{\alpha i(d-i/2+1/2)}\prod^{k}_{l=k-i+1}\big(1-e^{-l\alpha}\big)-\sum^{k}_{i=1}(-1)^{i+1}e^{\alpha i(d-i/2+1/2)}\bigg|$
$\leq\sum^{k}_{i=k/2+1}e^{\alpha i(d-i/2+1/2)}\left(1-\prod^{k}_{l=k-i+1}\big(1-e^{-l\alpha}\big)\right)+\sum^{k/2}_{i=1}e^{\alpha i(d-i/2+1/2)}\left(1-\prod^{k}_{l=k-i+1}\big(1-e^{-l\alpha}\big)\right).$ (S.7)
Suppose that $k>2d$. We bound the first sum of (S.7). By basic inequalities, we have
$\sum^{k}_{i=k/2+1}e^{\alpha i(d-i/2+1/2)}\left(1-\prod^{k}_{l=k-i+1}\big(1-e^{-l\alpha}\big)\right)\leq\sum^{k}_{i=k/2+1}e^{\alpha i(d-i/2+1/2)}\leq\frac{k}{2}\cdot e^{\alpha(k/2)(d-k/4+1/2)}.$
When $k\to\infty$, this final expression will go 0, so there exists some $K_{1}$ such that for all $k>K_{1}$, the final expression is less than $\epsilon/2$.
We perform a similar computation to bound the second sum in expression (S.7). Note that
$\prod^{k}_{l=k-i+1}\big(1-e^{-l\alpha}\big)\geq\big(1-e^{-(k-i+1)\alpha}\big)^{i}\geq\big(1-e^{-(k/2+1)\alpha}\big)^{k/2}.$
Thus,
$1-\prod^{k}_{l=k-i+1}\big(1-e^{-l\alpha}\big)\leq 1-(1-e^{-(k/2+1)\alpha})^{k/2}\leq\frac{k}{2}\cdot e^{-(k/2+1)\alpha}.$ (S.8)
Substituting expression (S.8) back into the second sum of expression (S.7), we have that
$\sum^{k/2}_{i=1}e^{\alpha i(d-i/2+1/2)}\left(1-\prod^{k}_{l=k-i+1}\big(1-e^{-l\alpha}\big)\right)$ $\leq\frac{k}{2}\sum^{k/2}_{i=1}e^{\alpha i(d-i/2+1/2)}e^{-(k/2+1)\alpha}$
$\leq\frac{k}{2}\cdot\frac{k}{2}\cdot e^{\alpha d(d-d/2+1/2)}e^{-(k/2+1)\alpha}.$
Note that as $k\to\infty$, this final expression goes to 0, so there exists some $K_{2}$ such that for all $k>K_{2}$, this expression is less than $\epsilon/2$.
Therefore, for all $k>\max(K_{1},K_{2})$,
$\bigg|\tilde{w}_{k-d}^{k,\gamma}-\sum^{k}_{i=1}(-1)^{i+1}e^{\alpha i(d-i/2+1/2)}\bigg|$
$\leq\frac{k}{2}\cdot e^{\alpha(k/2)(d-k/4+1/2)}+\frac{k^{2}}{4}\cdot e^{\alpha(d(d-d/2+1/2)-(k/2+1))}$
$\leq\epsilon/2+\epsilon/2=\epsilon.$
Thus, we have shown that $\tilde{w}_{j}^{k,\gamma}$ has a limit, and the limit is
$\check{w}_{d}^{\gamma}$ $=\lim\limits_{\begin{subarray}{c}k\to\infty,\,j\to\infty,\\ k-j=d\end{subarray}}\tilde{w}_{j}^{k,\gamma}$
$=\sum^{\infty}_{i=1}(-1)^{i+1}e^{\alpha i(d-i/2+1/2)}.\qed$
Now, we show that this form we just proved is the same as (11) to finish the proof of the property (a). Using the difference of squares formula,
$\sum^{\infty}_{i=1}(-1)^{i+1}e^{\alpha i(d-i/2+1/2)}$ $=\sum^{\infty}_{i=1}(-1)^{i+1}e^{(\alpha/2)[(d+1/2)^{2}-(i-(d+1/2))^{2}]}$
$=e^{(\alpha/2)(d+1/2)^{2}}\sum^{\infty}_{i=1}(-1)^{i+1}e^{-(\alpha/2)(i-(d+1/2))^{2}}.$
However, we can cancel some terms in the following way. For $d\geq 0$,
$\sum^{2d}_{i=1}(-1)^{i+1}e^{-(\alpha/2)(i-(d+1/2))^{2}}=\sum^{d}_{i=1}(-1)^{i+1}\big(e^{-(\alpha/2)(i-(d+1/2))^{2}}-e^{-(\alpha/2)((d+1/2)-i)^{2}}\big)=0.$ (S.9)
Similarly, for $d<0$,
$\sum^{2d+1}_{i=0}(-1)^{i+1}e^{-(\alpha/2)(i-(d+1/2))^{2}}=0.$ (S.10)
Therefore, using the cancellation of (S.9) and (S.10),
$e^{(\alpha/2)(d+1/2)^{2}}\sum^{\infty}_{i=1}(-1)^{i+1}e^{-(\alpha/2)(i-(d+1/2))^{2}}$
$\qquad=e^{(\alpha/2)(d+1/2)^{2}}\sum^{\infty}_{i=2d+1}(-1)^{i+1}e^{-(\alpha/2)(i-(d+1/2))^{2}}$
$\qquad=e^{(\alpha/2)(d+1/2)^{2}}\sum^{\infty}_{i=0}(-1)^{i+(2d+1)+1}e^{-(\alpha/2)(i+(2d+1)-(d+1/2))^{2}}$
$\qquad=\sum^{\infty}_{i=0}(-1)^{i}e^{-\alpha i(d+i/2+1/2)}.$
To prove property (b), using the closed form from (11), note that
$1-\check{w}_{-(d+1)}^{\gamma}$ $=1-\sum^{\infty}_{i=1}(-1)^{i+1}e^{\alpha i(-(d+1)-i/2+1/2)}$
$=1+\sum^{\infty}_{i=1}(-1)^{i}e^{-\alpha i(d+i/2+1/2)}$
$=\sum^{\infty}_{i=0}(-1)^{i}e^{-\alpha i(d+i/2+1/2)}$
$=\check{w}_{d}^{\gamma}.$
Therefore, $\check{w}_{d}^{\gamma}$ satisfies symmetry property $\check{w}_{d}^{\gamma}=1-\check{w}_{-(d+1)}^{\gamma}$.
To prove property (c), by basic asymptotic results, we write the bounds from (10) as
$\lim\limits_{\begin{subarray}{c}k\to\infty,\,j\to\infty,\\ k-j=d\end{subarray}}\frac{1-e^{-\alpha k}}{1-e^{-\alpha k}+e^{-\alpha(k-j)}}$ $=\frac{1}{1+e^{-\alpha d}},\text{ and}$
$\lim\limits_{\begin{subarray}{c}k\to\infty,\,j\to\infty,\\ k-j=d\end{subarray}}\frac{1-e^{-\alpha k}}{1-e^{-\alpha k}+e^{-\alpha(k-j+1)}}$ $=\frac{1}{1+e^{-\alpha(d+1)}}.$
Therefore, $\check{w}_{d}^{\gamma}$ satisfies
$\frac{1}{1+e^{-\alpha d}}\leq\check{w}_{d}^{\gamma}\leq\frac{1}{1+e^{-\alpha(d+1)}}.$ (S.11)
Additionally, using the bounds from (S.11), we have
$\check{w}_{d}^{\gamma}\leq\frac{1}{1+e^{-\alpha(d+1)}}\leq\check{w}_{d+1}^{\gamma},$ (S.12)
demonstrating that $\check{w}_{d}^{\gamma}$ is non-decreasing for $d\in\mathbb{Z}$. The inequality $\check{w}_{d}^{\gamma}\geq\frac{1}{1+e^{-\alpha(d+1/2)}}$ for $d\geq 0$ follows directly from (S.18). Furthermore, we obtain $\check{w}_{d}^{\gamma}\leq\frac{1}{1+e^{-\alpha(d+1/2)}}$ for $d<0$ by using the relation $\check{w}_{d}^{\gamma}=1-\check{w}_{-(d+1)}^{\gamma}$.
To prove property (d), note that property (i) in the footnote follows immediately from (S.12). Next, we prove properties (ii) and (iii) in the footnote. To this end, using the two different series representations (11) and (S.6) of $\check{w}_{d}^{\gamma}$ for integer $d$, we extend the definition of $\check{w}_{d}^{\gamma}$ to all real $d$ as follows. When d<1/2d<-1/2, let $\check{w}_{d}^{\gamma}=\sum_{i=1}^{\infty}(-1)^{i+1}e^{\alpha i(d-i/2+1/2)}$ and when $d\geq-1/2$, let $\check{w}_{d}^{\gamma}=\sum_{i=0}^{\infty}(-1)^{i}e^{-\alpha i(d+i/2+1/2)}$. For d>1/2d>-1/2, note that $\check{w}_{d}^{\gamma}$ has second derivative
$\frac{\partial^{2}}{\partial d^{2}}\check{w}_{d}^{\gamma}=\alpha^{2}\sum_{i=0}^{\infty}(-1)^{i}i^{2}e^{-\alpha i(d+i/2+1/2)}=\alpha^{2}\sum_{i=0}^{\infty}(-1)^{i}z^{2i}i^{2}q^{i^{2}},$ (S.13)
and $z,q\in(0,1)$, where $q=e^{-\alpha/2}$ and $z=e^{-(\alpha/2)(d+1/2)}$. Now, consider the following identity from (coogan2003q, Proposition 1.1):
$\sum_{i=0}^{\infty}(-1)^{i}z^{2i}q^{i^{2}}=\frac{1}{1+z}\sum_{i=0}^{\infty}z^{i}\prod_{j=0}^{i-1}\frac{1-zq^{j}}{1+zq^{j+1}},\quad|z|<1,\;|q|\leq 1.$ (S.14)
Differentiating both sides of (S.14) with respect to $q$ gives
$\sum_{i=0}^{\infty}(-1)^{i}z^{2i}i^{2}q^{i^{2}-1}=\frac{1}{1+z}\sum_{i=1}^{\infty}z^{i}\frac{\partial}{\partial q}\prod_{j=0}^{i-1}\frac{1-zq^{j}}{1+zq^{j+1}}.$ (S.15)
By the product rule and quotient rule on (S.15),
$\frac{\partial}{\partial q}\prod_{j=0}^{i-1}\frac{1-zq^{j}}{1+zq^{j+1}}=\sum_{l=0}^{i-1}\frac{q^{l-1}z(q^{l+1}z-l(q+1)-q)}{(1+zq^{l+1})^{2}}\prod_{0\leq j\leq i-1,\,j\neq l}\frac{1-zq^{j}}{1+zq^{j+1}}.$ (S.16)
Since $q^{l+1}z-l(q+1)-q\leq 0$ for all $l\geq 0$ and $(1-zq^{j})/(1+zq^{j+1})\geq 0$ for all $j\geq 0$ when $z,q\in[0,1)$, it follows that (S.16) is non-positive, implying the same for (S.15). Returning to (S.13), we deduce that $\check{w}_{d}^{\gamma}$ is concave for all d>1/2d>-1/2. An analogous argument shows that $\check{w}_{d}^{\gamma}$ is convex for d<1/2d<-1/2. This proves properties (ii) and (iii) in the footnote when $d\neq-1,0$. For the d=1,0d=-1,0 cases, we have from the just established convexity/concavity and continuity that
$\frac{1}{3}\check{w}_{-2}^{\gamma}+\frac{2}{3}\lim_{d\uparrow-1/2}\check{w}_{d}^{\gamma}$ $\geq\check{w}_{-1}^{\gamma},$ (S.17)
$\frac{2}{3}\check{w}_{-1/2}^{\gamma}+\frac{1}{3}\check{w}_{1}^{\gamma}$ $\leq\check{w}_{0}^{\gamma},$
where $\lim_{d\uparrow-1/2}\check{w}_{d}^{\gamma}=1-\check{w}_{-1/2}^{\gamma}$. We will show that these equations (S.17) together imply properties (ii) and (iii) in the footnote when d=1,0d=-1,0. Expanding the power series (S.14), ones obtains
$\check{w}_{d}^{\gamma}=\frac{1}{1+z}\sum_{i=0}^{\infty}z^{i}\prod_{j=0}^{i-1}\frac{1-zq^{j}}{1+zq^{j+1}}\geq\frac{1}{1+z}\sum_{i=0}^{\infty}z^{i}\prod_{j=0}^{i-1}\frac{1-z}{1+z}=\frac{1}{1+z^{2}}=\frac{1}{1+e^{-\alpha(d+1/2)}}.$ (S.18)
Thus, we have $\check{w}_{-1/2}^{\gamma}\geq 1/2$ and hence $\check{w}_{-1/2}^{\gamma}\geq\lim_{d\uparrow-1/2}\check{w}_{d}^{\gamma}$. Thus, from (S.17), we obtain
$3\check{w}_{-1}^{\gamma}+\check{w}_{1}^{\gamma}\leq 3\check{w}_{0}^{\gamma}+\check{w}_{-2}^{\gamma}.$ (S.19)
Finally, the identity $\check{w}_{d}^{\gamma}=1-\check{w}_{-(d+1)}^{\gamma}$ for $d\in\mathbb{Z}$ from part (c) of Theorem 3.2 applied to (S.19) yields properties (ii) and (iii) in the footnote when d=1,0d=-1,0. ∎
S.1.3 Proof of Theorem 4.1
From the decompositions (6) and (5), we have the following identity
$\|\mathbf{y}-\hat{\mathbf{f}}_{k,p}\|^{2}-\|\mathbf{y}-\hat{\mathbf{f}}_{k,m,p}\|^{2}=\sum_{j=k+1}^{p}\hat{\beta}^{2}_{j,p}\big(1-\big(-w_{j}^{k,m,p}\big)^{2}\big)-\sum_{j=1}^{k}\hat{\beta}^{2}_{j,p}\big(1-w_{j}^{k,m,p}\big)^{2}.$ (S.20)
Let $m_{p}=\lfloor\gamma p\rfloor$, where $0<\gamma<1$. Then, using (S.20), we have
$\mathbb{E}\Big[\|\mathbf{y}-\hat{\mathbf{f}}_{k,p}\|^{2}-\min_{m}\|\mathbf{y}-\hat{\mathbf{f}}_{k,m,p}\|^{2}\Big]$
$\qquad\geq\sum_{j=k+1}^{2k}\mathbb{E}[\hat{\beta}^{2}_{j,p}]\big(1-\big(-w_{j}^{k,m_{p},p}\big)^{2}\big)-\sum_{j=1}^{k}\mathbb{E}[\hat{\beta}^{2}_{j,p}]\big(1-w_{j}^{k,m_{p},p}\big)^{2}.$ (S.21)
To simplify further, we use Chebyshev’s sum inequality (hardy1988inequalities), noting that $\mathbb{E}[\hat{\beta}^{2}_{j,p}]$ is non-increasing in $j$ by definition and $w_{j}^{k,m_{p},p}$ is non-increasing in $j$ by Theorem 3.1. Thus, the second term in (S.21) is upper bounded by
$\sum_{j=1}^{k}\mathbb{E}[\hat{\beta}^{2}_{j,p}]\big(1-w_{j}^{k,m_{p},p}\big)^{2}$ $\leq\Bigg(\frac{1}{k}\sum_{j=1}^{k}\mathbb{E}[\hat{\beta}^{2}_{j,p}]\Bigg)\Bigg(\sum_{j=1}^{k}\big(1-w_{j}^{k,m_{p},p}\big)^{2}\Bigg),$ (S.22)
and the first term in (S.21) is lower bounded by
$\sum_{j=k+1}^{2k}\mathbb{E}[\hat{\beta}^{2}_{j,p}]\big(1-\big(1-w_{j}^{k,m_{p},p}\big)^{2}\big)$
$\qquad\qquad\geq\Bigg(\frac{1}{k}\sum_{j=k+1}^{2k}\mathbb{E}[\hat{\beta}^{2}_{j,p}]\Bigg)\Bigg(\sum_{j=k+1}^{2k}\big(1-\big(1-w_{j}^{k,m_{p},p}\big)^{2}\big)\Bigg).$ (S.23)
We claim that for a fixed $k$ and a sequence $\{\gamma_{p}\}$ that depends on both $p$ and $k$, with $\gamma_{p}\geq C>0$ where $C$ is some constant depending on $k$, if we set $m_{p}=\lfloor\gamma_{p}p\rfloor$, then there exists a constant $C(k,\epsilon)$ such that when $p(1-\gamma_{p})^{3}\geq C(k,\epsilon)$,
$\sum_{j=k+1}^{2k}\big(1-\big(-w_{j}^{k,m_{p},p}\big)^{2}\big)\geq(1-\epsilon)\sum_{j=k+1}^{2k}\big(1-\big(-\tilde{w}_{j}^{k,\gamma_{p}}\big)^{2}\big),$ (S.24)
and
$\sum_{j=1}^{k}\big(1-w_{j}^{k,m_{p},p}\big)^{2}\leq(1+\epsilon)\sum_{j=1}^{k}\big(1-\tilde{w}_{j}^{k,\gamma_{p}}\big)^{2}.$ (S.25)
For simplicity of notation, throughout the proof of this theorem, we use the notation $A\lesssim B$” to mean that there exists a constant $C$ depending on $k$ such that $A\leq CB$. To prove the claim, we first show that when $p(1-\gamma_{p})^{2}$ is sufficiently large,
$\frac{\Big|\sum_{j=k+1}^{2k}\big(1-\big(-w_{j}^{k,m_{p},p}\big)^{2}\big)-\sum_{j=k+1}^{2k}\big(1-\big(-\tilde{w}_{j}^{k,\gamma_{p}}\big)^{2}\big)\Big|}{\sum_{j=k+1}^{2k}\big(1-\big(-\tilde{w}_{j}^{k,\gamma_{p}}\big)^{2}\big)}\lesssim\frac{1}{p(1-\gamma_{p})^{2}}.$ (S.26)
If (S.26) holds, then it is clear that the claim for (S.24) holds.
To prove (S.26), we revisit the proof of Lemma S.1.1. Note that
$\frac{\binom{p-j}{m_{p}-1}}{\binom{p}{m_{p}}}=\frac{m_{p}}{p-j-m_{p}+1}\prod^{j}_{i=1}\frac{i+p-m_{p}}{i+p}=\frac{1+p-m_{p}}{p-j-m_{p}+1}\frac{m_{p}}{1+p}\prod^{j}_{i=2}\frac{i+p-m_{p}}{i+p}.$
For a fixed $2\leq j\leq 2k$, we have that for any $2\leq i\leq j$, $\big|\frac{i+p-m_{p}}{i+p}-(1-\gamma_{p})\big|\lesssim\frac{1}{p+i}$. Additionally, $|\frac{m_{p}}{1+p}-\gamma_{p}|\lesssim\frac{1}{p+1}$. Thus, we know when $p$ is sufficiently large, $\big|\frac{m}{1+p}\prod^{j}_{i=2}\frac{i+p-m}{i+p}-\gamma(1-\gamma)^{j-1}\big|\lesssim\frac{1}{p+1}$. Next, we observe that when $p(1-\gamma_{p})$ is sufficiently large,
$\frac{1+p-m_{p}}{p-j-m_{p}+1}-1=\frac{j}{p-j-m_{p}+1}\lesssim\frac{1}{p(1-\gamma_{p})}.$ (S.27)
Therefore, when $p(1-\gamma_{p})$ is sufficiently large, we have
$\left|\frac{\binom{p-j}{m_{p}-1}}{\binom{p}{m_{p}}}-\gamma_{p}(1-\gamma_{p})^{j-1}\right|\lesssim\frac{1}{p(1-\gamma_{p})}.$
Additionally, using a similar analysis, we know that when $p(1-\gamma_{p})$ is sufficiently large,
$\bigg|\frac{\binom{p-j}{m_{p}}}{\binom{p}{m_{p}}}-\big(1-\gamma_{p})^{j}\bigg|=\bigg|\prod_{i=0}^{j-1}\frac{p-m_{p}-i}{p-i}-\big(1-\gamma_{p})^{j}\bigg|\lesssim\frac{1}{p},$
and
$\left|\frac{\binom{p-j+1}{m_{p}}}{\binom{p}{m_{p}}}-\big(1-\gamma_{p})^{j-1}\right|\lesssim\frac{1}{p(1-\gamma_{p})}.$
According to the recurrence (8) and the fact that $w_{j}^{0,m_{p},p}=0$ for all $j$ and all $p$, we conclude that for any $1\leq j\leq 2k$, the quantity $w_{j}^{k,m_{p},p}$ can be determined through a finite number of recursive steps (depending on $k$). In other words, the approximation error $\big|w_{j}^{k,m_{p},p}-\tilde{w}_{j}^{k,\gamma_{p}}\big|$ accumulates only over a finite number of recursive steps (depending on $k$). Therefore, for any $1\leq j\leq 2k$, when $p(1-\gamma_{p})$ is sufficiently large, we have the error bound $\big|w_{j}^{k,m_{p},p}-\tilde{w}_{j}^{k,\gamma_{p}}\big|\lesssim\frac{1}{p(1-\gamma_{p})}$. Thus, it holds that
$\Bigg|\sum_{j=k+1}^{2k}\big(1-\big(-w_{j}^{k,m_{p},p}\big)^{2}\big)-\sum_{j=k+1}^{2k}\big(1-\big(-\tilde{w}_{j}^{k,\gamma_{p}}\big)^{2}\big)\Bigg|\lesssim\frac{1}{p(1-\gamma_{p})}.$ (S.28)
Next, we show that when $\gamma_{p}\gtrsim 1$, it holds that
$\sum_{j=k+1}^{2k}\big(1-\big(-\tilde{w}_{j}^{k,\gamma_{p}}\big)^{2}\big)\gtrsim 1-\gamma_{p}.$ (S.29)
To lower bound $\sum_{j=k+1}^{2k}\big(1-\big(-\tilde{w}_{j}^{k,\gamma_{p}}\big)^{2}\big)$ from (S.29), note that
$1-\big(1-\tilde{w}_{j}^{k,\gamma_{p}}\big)^{2}=\tilde{w}_{j}^{k,\gamma_{p}}\big(2-\tilde{w}_{j}^{k,\gamma_{p}}\big),$ (S.30)
so we wish to have simple upper and lower bounds on $\tilde{w}_{j}^{k,\gamma_{p}}$ for $j\geq k+1$ to lower bound each $1-\big(1-\tilde{w}_{j}^{k,\gamma_{p}}\big)^{2}$. Recall that $1-\gamma_{p}=e^{-\alpha_{p}}$. Since $j\geq k+1$, we have the following lower bound $\tilde{w}_{j}^{k,\gamma_{p}}$ from (10):
$\tilde{w}_{j}^{k,\gamma_{p}}\geq\frac{1-e^{-\alpha_{p}k}}{1-e^{-\alpha k}+e^{-\alpha_{p}(k-j)}}\geq e^{-\alpha_{p}(j-k)}\frac{1-e^{-\alpha_{p}k}}{1+e^{-\alpha_{p}}},\quad j\geq k+1.$ (S.31)
Now, from (10), observe that for all $j\geq k+1$,
$\tilde{w}_{j}^{k,\gamma_{p}}\leq\frac{1-e^{-\alpha_{p}k}}{1-e^{-\alpha_{p}k}+e^{-\alpha_{p}(k-j+1)}}\leq\frac{1-e^{-\alpha_{p}k}}{1-e^{-\alpha_{p}k}+1}\leq 1/2.$ (S.32)
From (S.31) and (S.32), we obtain
$1-(1-\tilde{w}_{j}^{k,\gamma_{p}}\big)^{2}=\tilde{w}_{j}^{k,\gamma_{p}}(2-\tilde{w}_{j}^{k,\gamma_{p}})\geq\frac{3}{2}e^{-\alpha_{p}(j-k)}\frac{1-e^{-\alpha_{p}k}}{1+e^{-\alpha_{p}}},\quad j\geq k+1.$ (S.33)
Substituting in (S.33), we thus have
$\sum_{j=k+1}^{2k}\big(1-\big(-\tilde{w}_{j}^{k,\gamma_{p}}\big)^{2}\big)\geq\frac{3}{2}e^{-\alpha_{p}}\frac{1-e^{-\alpha_{p}k}}{1-e^{-\alpha_{p}}}\frac{1-e^{-\alpha_{p}k}}{1+e^{-\alpha_{p}}}=\frac{3}{2}\frac{e^{-\alpha_{p}}(1-e^{-\alpha_{p}k})^{2}}{1-e^{-2\alpha_{p}}}.$ (S.34)
Noting that $\gamma_{p}\gtrsim$ implies $\alpha_{p}\gtrsim 1$, we obtain
$\sum_{j=k+1}^{2k}\big(1-\big(-\tilde{w}_{j}^{k,\gamma_{p}}\big)^{2}\big)\geq\frac{3}{2}e^{-\alpha_{p}}(1-e^{-\alpha_{p}k})^{2}\gtrsim e^{-\alpha_{p}}=1-\gamma_{p},$
which establishes (S.29). Combining (S.28) and (S.29) proves (S.26), which in turn confirms (S.24).
To prove (S.25), we show that when $\gamma_{p}\gtrsim 1$,
$\sum_{j=1}^{k}\big(1-\tilde{w}_{j}^{k,\gamma_{p}}\big)^{2}\gtrsim\big(1-\gamma_{p})^{2}.$ (S.35)
When $p(1-\gamma_{p})^{3}$ is sufficiently large, if (S.35) holds, then using the fact that $\big|w_{j}^{k,m_{p},p}-\tilde{w}_{j}^{k,\gamma_{p}}\big|\lesssim\frac{1}{p(1-\gamma_{p})}$, we know
$\frac{\Big|\sum_{j=1}^{k}\big(1-w_{j}^{k,m_{p},p}\big)^{2}-\sum_{j=1}^{k}\big(1-\tilde{w}_{j}^{k,\gamma_{p}}\big)^{2}\Big|}{\sum_{j=1}^{k}\big(1-\tilde{w}_{j}^{k,\gamma_{p}}\big)^{2}}\lesssim\frac{1}{p(1-\gamma_{p})^{3}},$
which implies (S.25).
Note that from (10), we have
$\tilde{w}_{j}^{k,\gamma_{p}}\leq\frac{1-e^{-\alpha_{p}k}}{1-e^{-\alpha_{p}k}+e^{-\alpha_{p}(k-j+1)}}\leq\frac{1}{1+e^{-\alpha_{p}(k-j+1)}}.$
Thus,
$\sum_{j=1}^{k}\big(1-\tilde{w}_{j}^{k,\gamma_{p}}\big)^{2}\geq\sum_{j=1}^{k}\frac{e^{-2\alpha_{p}(k-j+1)}}{\big(1+e^{-\alpha_{p}(k-j+1)}\big)^{2}}\geq\frac{1}{4}\sum_{j=1}^{k}e^{-2\alpha_{p}(k-j+1)}=\frac{e^{-2\alpha_{p}}(1-e^{-2\alpha_{p}k})}{4(1-e^{-2\alpha_{p}})}.$
Noting that $\gamma_{p}\gtrsim$ implies $\alpha_{p}\gtrsim 1$, we obtain (S.35).
Therefore, the claim is proven for (S.24) and (S.25). To upper bound the sum $\sum_{j=1}^{k}\big(1-\tilde{w}_{j}^{k,\gamma_{p}}\big)^{2}$ from (S.25), we use the following rearranged version of recurrence (S.4):
$1-\tilde{w}_{j}^{k,\gamma_{p}}=e^{-\alpha_{p}k}(1-\tilde{w}_{j-1}^{k,\gamma_{p}})+e^{-\alpha_{p}(k-j+1)}\tilde{w}_{j-1}^{k,\gamma_{p}}.$ (S.36)
Upon squaring both sides of (S.36), we obtain
$\big(1-\tilde{w}_{j}^{k,\gamma_{p}}\big)^{2}$ $=e^{-2\alpha_{p}k}\big(1-\tilde{w}_{j-1}^{k,\gamma_{p}}\big)^{2}+$
$\qquad 2e^{-\alpha_{p}(2k-j+1)}\tilde{w}_{j-1}^{k,\gamma_{p}}(1-\tilde{w}_{j-1}^{k,\gamma_{p}})+e^{-2\alpha_{p}(k-j+1)}\big(\tilde{w}_{j-1}^{k,\gamma_{p}}\big)^{2}.$ (S.37)
Summing (S.1.3) over $1\leq j\leq k$, using the Cauchy-Schwarz inequality, and the fact from Theorem 3.1 that $\tilde{w}_{j-1}^{k,\gamma_{p}}\geq\tilde{w}_{j}^{k,\gamma_{p}}$, we obtain
$\sum_{j=1}^{k}\big(1-\tilde{w}_{j}^{k,\gamma_{p}}\big)^{2}$ $\leq e^{-2\alpha_{p}k}\sum_{j=1}^{k}\big(1-\tilde{w}_{j-1}^{k,\gamma_{p}}\big)^{2}$
$\qquad+2e^{-\alpha_{p}k}\sqrt{\sum_{j=1}^{k}\big(1-\tilde{w}_{j-1}^{k,\gamma_{p}}\big)^{2}\sum_{j=1}^{k}e^{-2\alpha_{p}(k-j+1)}\big(\tilde{w}_{j-1}^{k,\gamma_{p}}\big)^{2}}$
$\qquad\qquad+\sum_{j=1}^{k}e^{-2\alpha_{p}(k-j+1)}\big(\tilde{w}_{j-1}^{k,\gamma_{p}}\big)^{2}$
$\leq e^{-2\alpha_{P}k}\sum_{j=1}^{k}\big(1-\tilde{w}_{j}^{k,\gamma_{p}}\big)^{2}$
$\qquad+2e^{-\alpha_{p}k}\sqrt{\sum_{j=1}^{k}\big(1-\tilde{w}_{j}^{k,\gamma_{p}}\big)^{2}\sum_{j=1}^{k}e^{-2\alpha_{p}(k-j+1)}\big(\tilde{w}_{j-1}^{k,\gamma_{p}}\big)^{2}}$ (S.38)
$\qquad\qquad+\sum_{j=1}^{k}e^{-2\alpha_{p}(k-j+1)}\big(\tilde{w}_{j-1}^{k,\gamma_{p}}\big)^{2}.$
Solving the quadratic inequality (S.1.3) yields
$\sum_{j=1}^{k}\big(1-\tilde{w}_{j}^{k,\gamma_{p}}\big)^{2}\leq\frac{\sum_{j=1}^{k}e^{-2\alpha_{p}(k-j+1)}\big(\tilde{w}_{j-1}^{k,\gamma_{p}}\big)^{2}}{(1-e^{-\alpha_{p}k})^{2}}.$ (S.39)
Now, from the upper bound (10), we have $\big(\tilde{w}_{j-1}^{k,\gamma_{p}}\big)^{2}\leq(1-e^{-\alpha_{p}k})^{2}$ for all $j\geq 1$. Substituting this into (S.39), we obtain the bound
$\sum_{j=1}^{k}\big(1-\tilde{w}_{j}^{k,\gamma_{p}}\big)^{2}\leq\sum_{j=1}^{k}e^{-2\alpha_{p}(k-j+1)}=\frac{e^{-2\alpha_{p}}(1-e^{-2\alpha_{p}k})}{1-e^{-2\alpha_{p}}}.$ (S.40)
To simplify notation, set
$A=\frac{1}{k}\sum_{j=1}^{k}\mathbb{E}[\hat{\beta}^{2}_{j,p}],\qquad B=\frac{1}{k}\sum_{j=k+1}^{2k}\mathbb{E}[\hat{\beta}^{2}_{j,p}],$ (S.41)
and note that $A\geq B$ by applying Definition 3.1. Using (S.41) and by the previous inequalities in (S.40) and (S.34), the training error gap (S.21) is at least
$\frac{3}{2}(1-\epsilon)B\frac{e^{-\alpha_{p}}(1-e^{-\alpha_{p}k})^{2}}{1-e^{-2\alpha_{p}}}-A(1+\epsilon)\frac{e^{-2\alpha_{p}}(1-e^{-2\alpha_{p}k})}{1-e^{-2\alpha_{p}}}$
$\qquad=\Big(\frac{1-e^{-\alpha_{p}k}}{1-e^{-\alpha_{p}}}\Big)\Big[\frac{3}{2}(1-\epsilon)B\frac{e^{-\alpha_{p}}(1-e^{-\alpha_{p}k})}{1+e^{-\alpha_{p}}}-(1+\epsilon)A\frac{e^{-2\alpha_{p}}(1+e^{-\alpha_{p}k})}{1+e^{-\alpha_{p}}}\Big]$
$\qquad\geq\frac{3}{2}(1-\epsilon)Be^{-\alpha_{p}}\frac{1-e^{-\alpha_{p}}}{1+e^{-\alpha_{p}}}-(1+\epsilon)Ae^{-2\alpha_{p}}$
$\qquad\geq\frac{3}{2}(1-\epsilon)Be^{-\alpha_{p}}\frac{1-e^{-\alpha_{0}}}{1+e^{-\alpha_{0}}}-(1+\epsilon)Ae^{-2\alpha_{p}},$ (S.42)
for any $\alpha_{0}\leq\alpha_{p}$. Take $\alpha_{p}^{\star}$ such that
$\alpha_{p}^{\star}=\log\Big(\frac{4}{3}\cdot\frac{1+e^{-\alpha_{0}^{\star}}}{1-e^{-\alpha_{0}^{\star}}}\cdot\frac{A}{B}\Big),$ (S.43)
where the corresponding $\gamma_{p}^{\star}$ is given by $\gamma_{p}^{\star}=1-e^{-\alpha_{p}^{\star}}$, and $\alpha_{0}^{\star}=(7+\sqrt{97})/6$. It is straightforward to verify that
$\alpha_{p}^{\star}=\log\Big(\frac{4}{3}\cdot\frac{1+e^{-\alpha_{0}^{\star}}}{1-e^{-\alpha_{0}^{\star}}}\cdot\frac{A}{B}\Big)\geq\log\Big(\frac{4}{3}\cdot\frac{1+e^{-\alpha_{0}^{\star}}}{1-e^{-\alpha_{0}^{\star}}}\Big)=\alpha_{0}^{\star}.$ (S.44)
Thus, substituting $\alpha_{p}^{\star}$ and $\alpha_{0}^{\star}$ into (S.42) and taking $\epsilon=0.05$ gives us the lower bound on expected training error gap (S.21) as
$\Big(\frac{3}{4}\Big)\Big(\frac{3}{4}-\frac{9}{4}\epsilon\Big)\Big(\frac{1-e^{-\alpha_{0}^{\star}}}{1+e^{-\alpha_{0}^{\star}}}\Big)^{2}\frac{\big(\frac{1}{k}\sum_{j=k+1}^{2k}\mathbb{E}[\hat{\beta}^{2}_{j,p}]\big)^{2}}{\frac{1}{k}\sum_{j=1}^{k}\mathbb{E}[\hat{\beta}^{2}_{j,p}]}$
$\qquad\geq\frac{1}{4}\frac{\big(\frac{1}{k}\sum_{j=k+1}^{2k}\mathbb{E}[\hat{\beta}^{2}_{j,p}]\big)^{2}}{\frac{1}{k}\sum_{j=1}^{k}\mathbb{E}[\hat{\beta}^{2}_{j,p}]}.$
Therefore, for any fixed $k\geq 1$, when $p(1-\gamma_{p}^{\star})^{3}\gtrsim 1$, we have
$\mathbb{E}\Big[\|\mathbf{y}-\hat{\mathbf{f}}_{k,p}\|^{2}-\min_{m}\|\mathbf{y}-\hat{\mathbf{f}}_{k,m,p}\|^{2}\Big]$ $\geq\frac{1}{4}\frac{\big(\frac{1}{k}\sum_{j=k+1}^{2k}\mathbb{E}[\hat{\beta}^{2}_{j,p}]\big)^{2}}{\frac{1}{k}\sum_{j=1}^{k}\mathbb{E}[\hat{\beta}^{2}_{j,p}]}.$
S.1.4 Proofs for Section 4.2
Proof of Theorem 4.2.
We claim that whenever $m^{\prime}>m$, the sequence $\big\{w_{j}^{k,m^{\prime},p}\big\}_{j=1}^{p}$ majorizes the sequence $\big\{w_{j}^{k,m,p}\big\}_{j=1}^{p}$. This means that for each $l\geq 1$,
$\sum_{j=1}^{l}w_{j}^{k,m^{\prime},p}\geq\sum_{j=1}^{l}w_{j}^{k,m,p},$ (S.45)
with equality when $l=p$.
Assuming the majorization in (S.45) holds, the theorem follows directly from Theorem A.3 in Chapter 5 from marshall11 using, in their notation, $g_{i}(z)=(\operatorname{df}(\hat{f}_{i,p})-\operatorname{df}(\hat{f}_{i-1,p}))z$.
We return to proving the claim. Let $S_{k,m,p}$ denote the indices of the features selected by one inner loop of EFS with parameters $k$, $m$, and $p$. We may write
$\begin{split}\sum_{j=1}^{l}w_{j}^{k,m,p}&=\sum_{j=1}^{l}\mathbb{P}(j\in S_{k,m,p})\\ &=\mathbb{E}\left[\sum_{j=1}^{l}\mathbb{1}(j\in S_{k,m,p})\right]\\ &=\mathbb{E}\left[Z_{k,m,p}(l)\right],\end{split}$ (S.46)
where $Z_{k,m,p}(l)=|S_{k,m,p}\cap[l]|$. Evidently, (S.45) will be established if we can produce a coupling between $S_{k,m,p}$ and $S_{k,m^{\prime},p}$ such that $Z_{k,m^{\prime}p}(l)\geq Z_{k,m,p}(l)$ for $l=1,2,\ldots,p$.
[…]
We now show the desired property inductively in $k$, with the base case $k=0$ trivial. Assume the inductive hypothesis. Let $b=\min B_{m}$. Then one can show that the index selected at iteration $k+1$ (with parameter $m$) is equal to $l_{*}=\min\left\{l\colon b\leq l-Z_{k,m,p}(l)\right\}$. We therefore have
$Z_{k+1,m,p}(l)=\begin{cases}Z_{k,m,p}(l)&\text{if}~l<l_{*}\\ Z_{k,m,p}(l)+1&\text{if}~l\geq l_{*}.\end{cases}$ (S.47)
Similarly, denoting $b^{\prime}=\min B_{m^{\prime}}$ and letting $l_{*}^{\prime}=\min\left\{l\colon b^{\prime}\leq l-Z_{k,m^{\prime},p}(l)\right\}$, we get
$Z_{k+1,m^{\prime},p}(l)=\begin{cases}Z_{k,m^{\prime},p}(l)&\text{if}~l<l_{*}^{\prime}\\ Z_{k,m^{\prime},p}(l)+1&\text{if}~l\geq l_{*}^{\prime}.\end{cases}$ (S.48)
To compare $Z_{k+1,m^{\prime},p}$ and $Z_{k,m,p}$, we further set $\tilde{l}=\min\left\{l\colon b\leq l-Z_{k,m^{\prime},p}(l)\right\}$ and construct the function $\tilde{Z}$ via
[…]
Since $b^{\prime}\leq b$, we have $\tilde{l}\geq l_{*}^{\prime}$, so that $\tilde{Z}\leq Z_{k+1,m^{\prime},p}$. Next, notice that
[…]
By the induction hypothesis, we have $Z_{k,m^{\prime},p}(l)-Z_{k,m,p}(l)\geq 0$. Meanwhile, on $l_{*}\leq l<\tilde{l}$, write
$Z_{k,m^{\prime},p}(l)-Z_{k,m,p}(l)=(Z_{k,m^{\prime},p}(l)-l)-(Z_{k,m,p}(l)-l).$ (S.51)
Because these functions take integral values, the only way $Z_{k,m^{\prime},p}(l)-Z_{k,m,p}(l)-1$ can be negative is if the right hand side of (S.51) is equal to 0. However, by the way in which $l_{*}$ is defined, we would have
$l-Z_{k,m^{\prime},p}(l)=l-Z_{k,m,p}(l)\geq b,$ (S.52)
which contradicts the definition of $\tilde{l}$. This implies that $Z_{k+1,m,p}\leq\tilde{Z}$, which when combined with the earlier comparison, gives $Z_{k+1,m,p}\leq Z_{k+1,m^{\prime},p}$ as desired. ∎
Utilizing the majorization in (S.45), we now turn to prove Theorem 5.1.
For the remaining proofs in this section, we depart from the notation used in the rest of this paper, and let $\mathbf{x}_{\cdot 1,p},\mathbf{x}_{\cdot 2,p},\ldots,\mathbf{x}_{\cdot p,p}$ denote a fixed ordering of the features. Correspondingly, their regression coefficients $\hat{\beta}_{1,p},\hat{\beta}_{2,p},\ldots,\hat{\beta}_{p,p}$ will no longer be assumed to be in sorted order.
Proof of Proposition 4.1.
Consider any pair of distinct feature indices, $i\neq j$. Then
$\hat{\beta}_{i,p}-\hat{\beta}_{j,p}=(\beta_{i,p}-\beta_{j,p})+\langle\mathbf{x}_{\cdot i,p}-\mathbf{x}_{\cdot j,p},\boldsymbol{\varepsilon}\rangle.$ (S.53)
The second term has the distribution $\mathcal{N}(0,2\sigma^{2}/n)$. In particular, if $i\leq k$ and $j>k$, we have
$\mathbb{P}\left(\hat{\beta}_{i}-\hat{\beta}_{j}<0\right)\leq\exp(-n/4\sigma^{2}).$ (S.54)
Similarly, we have
$\mathbb{P}\left(\hat{\beta}_{i}+\hat{\beta}_{j}<0\right)\leq\exp(-n/4\sigma^{2}).$ (S.55)
Putting these together gives
$\begin{split}\mathbb{P}\left(|\hat{\beta}_{j,p}|>|\hat{\beta}_{i,p}|\right)&\leq\mathbb{P}\left(|\hat{\beta}_{j,p}|>\hat{\beta}_{i,p}\right)\\ &\leq\mathbb{P}\left(\hat{\beta}_{j,p}>\hat{\beta}_{i,p}\right)+\mathbb{P}\left(-\hat{\beta}_{j,p}<\hat{\beta}_{i,p}\right)\\ &\leq 2\exp(-n/4\sigma^{2}).\end{split}$ (S.56)
Let $\mathcal{A}$ be the event on which $\operatorname{rank}(|\hat{\beta}_{j,p}|)\leq k$, for $j=1,\ldots,k$. We have
$\begin{split}\mathbb{P}\left(\mathcal{A}\right)&=\mathbb{P}\left(|\hat{\beta}_{i,p}|>|\hat{\beta}_{j,p}|,1\leq i\leq k<j\leq p\right)\\ &\geq 1-2k(p-k)\exp(-n/4\sigma^{2}).\end{split}$ (S.57)
Next, let $\widetilde{\mathbf{H}}_{k}=n^{-1}\sum_{i=1}^{k}\mathbf{x}_{\cdot i,k}\mathbf{x}_{\cdot i,k}^{\top}$. This is the projection matrix onto the column span of the first $k$ features. Let $\mathbf{H}_{k}$ be the smoothing matrix corresponding to $k$ steps of FS, noting that it depends on $\boldsymbol{\varepsilon}=(\epsilon_{1},\epsilon_{2},\ldots,\epsilon_{n})$. Observe that on the event $\mathcal{A}$, we have $\widetilde{\mathbf{H}}_{k}=\mathbf{H}_{k}$. We can then write
$\begin{split}\operatorname{df}(\hat{f}_{k,p})&=\frac{1}{\sigma^{2}}\mathbb{E}\left[\boldsymbol{\varepsilon}^{\top}\mathbf{H}_{k}\mathbf{y}\right]\\ &=\frac{1}{\sigma^{2}}\mathbb{E}\left[\boldsymbol{\varepsilon}^{\top}\mathbf{H}_{k}\mathbf{y}\cdot\mathbb{1}(\mathcal{A})\right]+\frac{1}{\sigma^{2}}\mathbb{E}\left[\boldsymbol{\varepsilon}^{\top}\mathbf{H}_{k}\mathbf{y}\cdot\mathbb{1}(\mathcal{A}^{c})\right]\\ &=\frac{1}{\sigma^{2}}\mathbb{E}\left[\boldsymbol{\varepsilon}^{\top}\widetilde{\mathbf{H}}_{k}\mathbf{y}\right]-\frac{1}{\sigma^{2}}\mathbb{E}\left[\boldsymbol{\varepsilon}^{\top}\widetilde{\mathbf{H}}_{k}\mathbf{y}\cdot\mathbb{1}(\mathcal{A}^{c})\right]+\frac{1}{\sigma^{2}}\mathbb{E}\left[\boldsymbol{\varepsilon}^{\top}\mathbf{H}_{k}\mathbf{y}\cdot\mathbb{1}(\mathcal{A}^{c})\right].\end{split}$ (S.58)
The first term satisfies
$\frac{1}{\sigma^{2}}\mathbb{E}\left[\boldsymbol{\varepsilon}^{\top}\widetilde{\mathbf{H}}_{k}\mathbf{y}\right]=\frac{1}{\sigma^{2}}\mathbb{E}\left[\boldsymbol{\varepsilon}^{\top}\widetilde{\mathbf{H}}_{k}\boldsymbol{\varepsilon}\right]=k.$ (S.59)
To bound the third term, notice that since $\mathbf{H}_{k}$ is a projection matrix, we can bound
$\begin{split}\boldsymbol{\varepsilon}^{\top}\mathbf{H}_{k}\mathbf{y}&=\boldsymbol{\varepsilon}^{\top}\mathbf{H}_{k}\boldsymbol{\varepsilon}+\boldsymbol{\varepsilon}^{\top}\mathbf{H}_{k}\mathbf{f}\\ &\leq\|\boldsymbol{\varepsilon}\|^{2}_{2}+\|\boldsymbol{\varepsilon}\|_{2}\|\mathbf{f}\|_{2}.\end{split}$ (S.60)
Using the fact that $\|\mathbf{f}\|_{2}^{2}=kn$, we get
$\mathbb{E}\left[(\boldsymbol{\varepsilon}^{\top}\mathbf{H}_{k}\mathbf{y})^{2}\right]^{1/2}\leq C(\sqrt{k}+\sigma)\sigma n,$ (S.61)
where $C$ is a universal constant. Applying Cauchy-Schwarz thus gives us
$\begin{split}\frac{1}{\sigma^{2}}\mathbb{E}\left[\boldsymbol{\varepsilon}^{\top}\mathbf{H}\mathbf{y}\cdot\mathbb{1}(\mathcal{A}^{c})\right]&\leq\frac{1}{\sigma^{2}}\mathbb{E}\left[(\boldsymbol{\varepsilon}^{\top}\mathbf{H}\mathbf{y})^{2}\right]^{1/2}\mathbb{P}\left(\mathcal{A}\right)^{1/2}\\ &\leq C(\sqrt{k}/\sigma+1)nk(p-k)\exp(-n/8\sigma^{2}).\end{split}$ (S.62)
By a similar calculation, the second term is bounded by the same value. Hence,
$\operatorname{df}(\hat{f}_{k,p})=k+C(\sqrt{k}/\sigma+1)nk(p-k)\exp(-n/8\sigma^{2}).$ (S.63)
Meanwhile, let $\tau$ denote the index of the feature selected on the $(k+1)$-th step of FS. We then have $\hat{f}_{k+1,p}-\hat{f}_{k,p}=\hat{\beta}_{\tau,p}x_{\cdot\tau,p}$, which gives
$\operatorname{df}(\hat{f}_{k+1,p})-\operatorname{df}(\hat{f}_{k,p})=\operatorname{df}(\hat{\beta}_{\tau}\psi_{\tau}).$ (S.64)
On the event $\mathcal{A}$, it is easy to see that
$\tau=\operatorname*{arg\,max}_{k+1\leq j\leq p}\hat{\beta}_{j,p}^{2}.$ (S.65)
As such, we compute
$\begin{split}\operatorname{df}(\hat{\beta}_{\tau}\psi_{\tau})&=\frac{1}{\sigma^{2}}\mathbb{E}\left[(\mathbf{y}-\mathbb{E}y)^{\top}\hat{\beta}_{\tau,p}\mathbf{x}_{\cdot j,p}\right]\\ &=\frac{n}{\sigma^{2}}\mathbb{E}\left[\hat{\beta}_{\tau,p}^{2}\right]\\ &\geq\mathbb{E}\left[\operatorname*{arg\,max}_{k+1\leq j\leq p}\frac{n}{\sigma^{2}}\hat{\beta}_{j,p}^{2}\mathbb{1}(\mathcal{A})\right]\\ &=\mathbb{E}\left[\operatorname*{arg\,max}_{k+1\leq j\leq p}\frac{n}{\sigma^{2}}\hat{\beta}_{j,p}^{2}\right]-\mathbb{E}\left[\operatorname*{arg\,max}_{k+1\leq j\leq p}\frac{n}{\sigma^{2}}\hat{\beta}_{j,p}^{2}\mathbb{1}(\mathcal{A}^{c})\right].\end{split}$ (S.66)
Since $n\hat{\beta}_{j,p}^{2}/\sigma^{2}\sim_{\text{IID}}\chi^{2}_{1}$ for $k+1\leq j\leq p$, Lemma S.1.6 implies that the first term is lower bounded by $\frac{1}{2}\log(p-k)$ for pkp-k large enough. For the second term, notice that
$\operatorname*{arg\,max}_{k+1\leq j\leq p}\frac{n}{\sigma^{2}}\hat{\beta}_{j,p}^{2}\leq\sum_{k+1\leq j\leq p}\frac{n}{\sigma^{2}}\hat{\beta}_{j,p}^{2}=\frac{1}{\sigma^{2}}\boldsymbol{\varepsilon}^{\top}\left(\sum_{j=k+1}^{p}\mathbf{x}_{\cdot j,p}\mathbf{x}_{\cdot j,p}^{\top}\right)\boldsymbol{\varepsilon}\leq\frac{\|\boldsymbol{\varepsilon}\|_{2}^{2}}{\sigma^{2}}.$ (S.67)
Applying Cauchy-Schwarz to the second term, we can bound it (in absolute value) from above by
$Cn\exp(-n/8\sigma^{2}).$ (S.68)
Hence,
$\operatorname{df}(\hat{f}_{k+1,p})-\operatorname{df}(\hat{f}_{k,p})\geq\frac{1}{2}\log(p-k)+Cn\exp(-n/8\sigma^{2}).$ (S.69)
Using (13), we get
$\operatorname{df}(\hat{f}_{k,m,p})\geq(w_{k}^{k,m,p}-w_{k+1}^{k,m,p})\operatorname{df}(\hat{f}_{k+1,p}).$ (S.70)
Putting everything together, we therefore get
$\operatorname{df}(\hat{f}_{k,m,p})-\operatorname{df}(\hat{f}_{k,p})\geq\frac{1}{2}(w_{k}^{k,m,p}-w_{k+1}^{k,m,p})\log(p-k)-k+C(\sqrt{k}/\sigma+1)nk(p-k)\exp(-n/8\sigma^{2}).$ (S.71)
Noting that Assumption 3.1 implies $k\leq p\leq n$ and taking a limit as $n,m,p$ go to infinity finishes the proof. ∎
Lemma S.1.6.
Let $Z_{1},Z_{2},\ldots,Z_{N}\sim_{\text{IID}}\chi^{2}_{1}$. For $N$ large enough, we have
$\mathbb{E}\left[\max_{1\leq i\leq N}Z_{i}\right]\geq\frac{1}{2}\log N.$ (S.72)
Proof.
$\begin{split}\mathbb{E}\left[\max_{1\leq i\leq N}Z_{i}\right]&=\int_{0}^{\infty}\mathbb{P}\left(\max_{1\leq i\leq N}Z_{i}\geq t\right)dt\\ &=\int_{0}^{\infty}1-\mathbb{P}\left(\max_{1\leq i\leq N}Z_{i}<t\right)dt\\ &=\int_{0}^{\infty}1-\mathbb{P}\left(Z_{1}<t\right)^{N}dt\\ &=\int_{0}^{\infty}1-(2\Phi(\sqrt{t})-1)^{N}dt,\end{split}$ (S.73)
where $\Phi(z)$ is the CDF of a Gaussian distribution. Using the Mills’ ratio bound, we have $\Phi(z)\leq 1-e^{-z^{2}}$ for $z$ large enough, which would imply that $(2\Phi(z)-1)^{N}\leq(1-2e^{-z^{2}})^{N}$. Since the integrand on the right hand side of (S.73) is decreasing, we have, for $N$ large enough, that
$\begin{split}\min_{0\leq t\leq\log N}1-(2\Phi(\sqrt{t})-1)^{N}&\geq 1-(2\Phi(\sqrt{\log N})-1)^{N}\\ &\geq 1-(1-2/N)^{N}\\ &\geq\frac{1}{2}.\end{split}$ (S.74)
Plugging this lower bound into (S.73) and integrating from 0 to $N$ completes the proof. ∎
S.1.5 Proof of Proposition 5.1
According to the decomposition (6), it suffices to prove
$\sum_{j=1}^{p}\hat{\beta}^{2}_{j,p}\big(w_{j}^{k,m,p}\big)^{2}\leq\sum_{j=1}^{p}\hat{\beta}^{2}_{j,p}\big(w_{j}^{k,m^{\prime},p}\big)^{2}.$ (S.75)
This follows immediately from Theorem A.3 in Chapter 5 from marshall11, where, in their notation, we take $g_{i}(z)=\beta^{2}_{i,p}z^{2}$.
S.1.6 Proof of Proposition 6.1
Denote the features indices selected by FS($p$) (in order) as $\pi(1),\pi(2),\ldots,\pi(p)$. Since
$\langle\mathbf{x}_{\cdot i,p},\mathbf{y}\rangle=\begin{cases}\beta&i=1,2\\ \zeta&3\leq i\leq p-1\\ 2\cdot 3^{-1/2}\beta&i=p,\end{cases}$ (S.76)
we have $\pi(1)=p$. Next, the first step residual satisfies
$\langle\mathbf{x}_{\cdot i,p},\mathbf{r}_{1}\rangle=\begin{cases}\beta/3&i=1,2\\ \zeta&3\leq i\leq p-1.\end{cases}$ (S.77)
Furthermore,
$\begin{split}\|\mathbf{P}_{\mathcal{M}_{1}}^{\perp}\mathbf{x}_{\cdot i,p}\|^{2}&=1-\langle\mathbf{x}_{\cdot i,p},\mathbf{x}_{\cdot p,p}\rangle^{2}\\ &=\begin{cases}2/3&i=1,2\\ 1&3\leq i\leq p-1.\end{cases}\end{split}$ (S.78)
Putting these together, the objective for each feature has the value
$\frac{|\langle\mathbf{x}_{\cdot i,p},\mathbf{r}_{1}\rangle|}{\|\mathbf{P}_{\mathcal{M}_{1}}^{\perp}\mathbf{x}_{\cdot i,p}\|}=\begin{cases}6^{-1/2}\beta&i=1,2\\ \zeta&3\leq i\leq p-1.\end{cases}$ (S.79)
As such, we have $\pi(k)=k-1$ for $k=2,3,\ldots,p$. This means that for any $k\geq 3$,
$\|\mathbf{y}-\hat{\mathbf{f}}_{k,p}\|^{2}=(p-k)\zeta^{2}.$ (S.80)
We now study the base estimator of EFS, denoted $\check{f}_{k,m,p}$, which depends on a random seed $\theta$. At iteration $l$, suppose $p\notin\mathcal{M}_{l}$. Observe that the residual $r_{l}$ satisfies
$\frac{|\langle\mathbf{x}_{\cdot p,p},\mathbf{r}_{l}\rangle|}{\|\mathbf{P}_{\mathcal{M}_{1}}^{\perp}\mathbf{x}_{\cdot i,p}\|}=\begin{cases}2\cdot 3^{-1/2}\beta&\text{if}~|\{1,2\}\cap\mathcal{M}_{l}|=0\\ 2^{-1/2}\beta&\text{if}~|\{1,2\}\cap\mathcal{M}_{l}|=1\\ 0&\text{if}~|\{1,2\}\cap\mathcal{M}_{l}|=2.\end{cases}$ (S.81)
Hence, if both $x_{1,p}$ and $x_{2,p}$ are selected before $x_{p,p}$, then $x_{p,p}$ is never selected. Denote the feature indices selected by EFS (in order) as $\tau(1),\tau(2),\ldots,\tau(p)$. Then $\tau^{-1}(j)$ denotes the step at which feature $x_{j,p}$ was selected. Note that these depend on the random seed $\theta$. We now decompose the probability space into three events. Let
$\mathcal{E}_{1}=\{\tau^{-1}(1),\tau^{-1}(2)\leq\min(\tau^{-1}(p),k)\},$ (S.82)
$\mathcal{E}_{2}=\{\tau^{-1}(1),\tau^{-1}(2)\leq k\}\backslash\mathcal{E}_{1},$ (S.83)
$\mathcal{E}_{3}=\left(\mathcal{E}_{1}\cup\mathcal{E}_{2}\right)^{c}.$ (S.84)
We have
$\|\mathbf{y}-\check{\mathbf{f}}_{k,m,p}\|^{2}\leq\begin{cases}(p-k-1)\zeta^{2}&\text{if}~\mathcal{E}_{1}~\text{holds},\\ (p-k)\zeta^{2}&\text{if}~\mathcal{E}_{2}~\text{holds},\\ (p-k)\zeta^{2}+2\beta^{2}&\text{if}~\mathcal{E}_{3}~\text{holds}.\end{cases}$ (S.85)
We next compute bounds for the event probabilities. For $l=1,\ldots,k$, let $S_{l}$ denote the set of $m$ indices sampled at iteration $l$. Denote $\gamma=m/p$ as before. By our previous discussion, we have
$\begin{split}\mathbb{P}(\mathcal{E}_{1})&\geq\mathbb{P}(1\in S_{1},0\notin S_{1},2\in S_{2},0\notin S_{2})\\ &=\frac{\binom{p-2}{m-1}\binom{p-3}{m-1}}{\binom{p}{m}\binom{p-1}{m}}\\ &=\gamma^{2}(1-\gamma^{2})+O(1/p).\end{split}$ (S.86)
To compute a bound for $\mathbb{P}(\mathcal{E}_{3})$, let us redefine the way $S_{l}$ is generated as follows: First draw a subset $\tilde{S}_{l}$ of size $m$ from $[p]$, i.e. including indices that have already been selected. If it contains an index that has already been selected, throw it out and draw it again. Let $W_{l}^{(j)}=\mathbb{1}(j\in\tilde{S}_{l})$ for $j=1,2$, $l=1,\ldots,k$. Then these are Bernoulli random features with probability $\gamma$. Note also that if $\sum_{l=1}^{k}W_{l}^{(j)}\geq 3$, then $\tau^{-1}(j)\leq k$. We thus have
$\begin{split}\mathbb{P}(\mathcal{E}_{3})&=1-\mathbb{P}(\tau^{-1}(1),\tau^{-1}(2)\leq k)\\ &=\mathbb{P}(\tau^{-1}(1)>k~\text{or}~\tau^{-1}(2)>k)\\ &\leq\mathbb{P}(\tau^{-1}(1)>k)+\mathbb{P}(\tau^{-1}(2)>k)\\ &\leq\mathbb{P}\left(\sum_{l=1}^{k}W_{l}^{(1)}<3\right)+\mathbb{P}\left(\sum_{l=1}^{k}W_{l}^{(2)}<3\right)\\ &\leq 2e^{-k\gamma}(ek\gamma/2)^{2},\end{split}$ (S.87)
where the last inequality follows from Chernoff’s inequality.
Putting everything together, and using Jensen’s inequality, we compute
$\begin{split}\mathbb{E}\big[\|\mathbf{y}-\hat{\mathbf{f}}_{k,p}\|^{2}\big]-\mathbb{E}\big[\|\mathbf{y}-\check{\mathbf{f}}_{k,m,p}\|^{2}\big]&\geq\zeta^{2}\mathbb{P}(\mathcal{E}_{1})-2\beta^{2}\mathbb{P}(\mathcal{E}_{3})\\ &\geq\zeta^{2}(1-\gamma^{2})\gamma^{2}-4\beta^{2}e^{-k\gamma}(ek\gamma/2)^{2}-O(1/p).\end{split}$ (S.88)
For $k$ and $p$ large enough, this last quantity is strictly positive.